### v1

# Diagonal Simplicial Tensor Modules and Algebraic $n$-Hypergroupoids

Abstract
Let $A$ be a commutative ring, $k\in\mathbb{Z}^{+}$, and $\vec{s}=(n_{1},\dots,n_{k})\in(\mathbb{Z}^{+})^{k}$ with $n=\min(\vec{s})-1$. We attach to $\vec{s}$ a diagonal simplicial tensor module $X_{\bullet}(\vec{s};A)$ whose $p$-simplices are functions on a cosimplicial index set $I_{p}(\vec{s})\subseteq\mathbb{N}^{k}$. This extends Quillen’s diagonal on double simplicial groups to $k$ commuting directions in $\mathbf{Mod}_{A}$, compatibly with Dold–Kan normalization.
We analyze the horn kernels $R_{p,j}(X)$ via “missing indices” and show that $R_{p,j}(X)\neq 0$ if and only if $k\geq p$, independently of $j$. It follows that $X_{\bullet}(\vec{s};A)$ is an algebraic $n$-hypergroupoid in the sense of Duskin–Glenn precisely when $k\leq n$, and that horn fillers in dimension $n$ are non-unique precisely when $k\geq n$; in particular $X_{\bullet}(\vec{s};A)$ is strict in our sense precisely when $k=n$. A Horn Non-Degeneracy Lemma then shows that $R_{n,j}(X)\cap D_{n}=\{0\}$ and yields a decomposition $X_{n}=R_{n,j}(X)\oplus D_{n}$. A shift-and-truncate chain homotopy, $\operatorname{Stab}(\vec{s})$-equivariant and filtration-preserving, shows that $X_{\bullet}(\vec{s};A)$ is contractible and forces the associated spectral sequence to collapse at $E_{1}$.
When $A$ is an infinite field $K$, we classify simplicial submodules generated by a single tensor via kernel sequences and a moduli map to a product of Grassmannians, obtaining an irreducible and unirational incidence variety.
keywords:
diagonal simplicial tensor modules , horn kernels , strict algebraic $n$-hypergroupoids , incidence varieties , moduli spaces
2020 MSC:
Primary 18G30; Secondary 55U10, 14M15, 05E45, 15A69.
[…]
[cuny]organization=The City University of New York, state=NY, country=USA
1 Introduction
Let $A$ be a commutative ring and $k\in\mathbb{Z}^{+}$. Fix a shape $\vec{s}=(n_{1},\ldots,n_{k})\in(\mathbb{Z}^{+})^{k}$ and set $n=\min(\vec{s})-1$. We construct the diagonal simplicial tensor module $X_{\bullet}(\vec{s};A)$ (DSTM), a simplicial $A$-module whose $p$-simplices are functions on cosimplicial index sets $I_{p}(\vec{s})\subseteq\mathbb{N}^{k}$.
This construction generalizes Quillen’s diagonal on a double simplicial group [Quillen1966] to $k$ commuting simplicial directions in the category $\mathbf{Mod}_{A}$ and is compatible with Dold–Kan normalization in the sense of Dold, Kan, May, Weibel, and Goerss–Jardine [Dold1958, Kan1958, MaySimplicial, Weibel, GoerssJardine].
In particular, the diagonal simplicial module $X_{\bullet}(\vec{s};A)$ is obtained from $X_{\bullet}(\vec{s};\mathbb{Z})$ by base change, and the normalized Moore complexes satisfy
$N_{\bullet}\bigl(X_{\bullet}(\vec{s};A)\bigr)\cong N_{\bullet}\bigl(X_{\bullet}(\vec{s};\mathbb{Z})\bigr)\otimes_{\mathbb{Z}}A$
for all commutative rings $A$.
We analyze the horn kernels
$R_{p,j}(X)=\bigcap_{i\neq j}\ker\bigl(d_{i}:X_{p}\to X_{p-1}\bigr),$
and show that their combinatorics controls the homotopy-theoretic behavior of the DSTM. For the simplicial dimension $n=\min(\vec{s})-1$ we prove that $X_{\bullet}(\vec{s};A)$ is a strict algebraic $n$-hypergroupoid in this sense if and only if the tensor order satisfies $k=n$. The transition from non-unique to unique horn fillers occurs exactly at $p=n$, and this threshold is determined by the tensor order.
We briefly outline the main aspects of the paper: combinatorics and classification of the DSTM, and the algebraic geometry of generated submodules.
Combinatorics and Classification
In the abelian setting of simplicial modules, horn fillers always exist (the Kan condition is automatic). The obstruction to the uniqueness of fillers is captured by the horn kernel $R_{p,j}$. Following terminology from Duskin–Glenn [Duskin1979, Glenn1982] and the nnLab, a simplicial module is a strict algebraic $n$-hypergroupoid if fillers are unique in dimensions strictly greater than $n$ ($R_{p,j}=0$ for $p>n$), and non-unique in dimension $n$ ($R_{n,j}\neq 0$).
[…]
Algebraic Geometry of Generated Submodules
When $A$ is an infinite field $K$, we work throughout with submodules generated by a single tensor $T\in X_{n}(\vec{s};K)$; we do not attempt to parametrize arbitrary subcomplexes. Such a generated submodule $\langle T\rangle$ is encoded by its kernel sequence
$K(T)_{\bullet}:=(\ker f_{T,p})_{p}\subseteq K[\Delta^{n}]_{\bullet},$
where $f_{T}$ is the realization map from Section 8.
We define a moduli map $\Psi$ from a Zariski-open locus $\mathcal{U}\subset X_{n}(\vec{s};K)$ to a product of Grassmannians by sending $T$ to the collection of subspaces $K(T)_{p}\subseteq K[\Delta^{n}]_{p}$. The image $\mathcal{M}(\vec{s})$ lies within a closed incidence subvariety defined by simplicial compatibility conditions, which we show are linear in the Segre–Plücker coordinates. This reduces the classification of these generated submodules to the geometry of these incidence conditions and of the index collision maps $\mathcal{I}_{p}$.
Organization
Section 2 establishes notation and the DSTM construction. Section 3 characterizes missing indices and the horn kernel basis. Section 4 develops the Moore filler algorithm and normalization. Section 5 establishes rank formulas and the classification theorem. Section 6 proves the Horn Non-Degeneracy Lemma and the decomposition $X_{n}=R_{n,j}\oplus D_{n}$. Section 7 develops the homology dichotomy and the hypergroupoid classification. Section 8 develops the algebraic geometry of generated submodules. A proves contractibility via an explicit equivariant homotopy.
2 Diagonal simplicial tensor modules
Throughout, let $A$ be a commutative ring and write $\mathbf{Mod}_{A}$ for the category of $A$-modules. Denote by $\mathbb{N}$ the set of nonnegative integers and by $\mathbb{Z}^{+}$ the set of positive integers.
2.1 Simplicial preliminaries
The simplicial category $\Delta$ has objects $p]:=\{0,\ldots,p\}$ for $p\in\mathbb{N}$; morphisms are nondecreasing maps. It is generated by the coface maps $\delta_{i}^{p}:[p-1]\to[p]$ (the injection missing $i$) and the codegeneracy maps $\sigma_{i}^{p}:[p+1]\to[p]$ (the surjection repeating $i$).
A simplicial $A$-module is a functor $X_{\bullet}:\Delta^{\mathrm{op}}\to\mathbf{Mod}_{A}$. We write $X_{p}:=X([p])$. The induced maps are the face maps $d_{i}^{p}:=X(\delta_{i}^{p}):X_{p}\to X_{p-1}$ and the degeneracy maps $s_{i}^{p}:=X(\sigma_{i}^{p}):X_{p}\to X_{p+1}$. The associated chain complex $(X_{\bullet},\partial_{\bullet})$ has boundary maps $\partial_{p}=\sum_{i=0}^{p}(-1)^{i}d_{i}^{p}$.
2.2 Setup and index sets
We now define the index sets underlying the diagonal simplicial tensor module. Let $k\in\mathbb{Z}^{+}$. We analyze modules consisting of tensors of order $k$. A shape is a $k$-tuple $\vec{s}=(n_{1},\dots,n_{k})\in(\mathbb{Z}^{+})^{k}$.
We fix the shape $\vec{s}$ throughout and define its simplicial dimension as
$n=n(\vec{s}):=\min(n_{1},\dots,n_{k})-1.$
For any degree $p\geq 0$ and axis $1\leq a\leq k$, we define the index bounds:
$M_{a}(p):=n_{a}-1-n+p.$
Since $n_{a}-1\geq n$, we have $M_{a}(p)\geq p$ for all $a$. The index set in degree $p$ is defined as
$I_{p}:=\prod_{a=1}^{k}[\,M_{a}(p)\,].$
Since $p]\subseteq[\,M_{a}(p)\,]$ for every $a$, we have the inclusion $p]^{k}\subseteq I_{p}$.
2.3 The Cosimplicial Index Set $I_{\bullet}$
We define a functor $I_{\bullet}:\Delta\to\mathbf{Set}$. On objects, $I_{\bullet}([p]):=I_{p}$. Note that $M_{a}(p\pm 1)=M_{a}(p)\pm 1$.
The structure maps are defined ”diagonally” as products of the standard generators.
Definition 2.1 (Cosimplicial Structure Maps).
The functor $I_{\bullet}$ maps the generators of $\Delta$ as follows:
For $p\geq 1$ and $0\leq i\leq p$, the index coface map $\Delta_{i}^{p}:I_{p-1}\to I_{p}$ is
$\Delta_{i}^{p}:=I_{\bullet}(\delta_{i}^{p}):=\prod_{a=1}^{k}\delta_{i}^{M_{a}(p)}.$
For $p\geq 0$ and $0\leq i\leq p$, the index codegeneracy map $\Sigma_{i}^{p}:I_{p+1}\to I_{p}$ is
$\Sigma_{i}^{p}:=I_{\bullet}(\sigma_{i}^{p}):=\prod_{a=1}^{k}\sigma_{i}^{M_{a}(p)}.$
Since $M_{a}(p)\geq p$, all factors are defined for the common index $i$. Because the standard maps $\delta_{i},\sigma_{i}$ satisfy the cosimplicial identities, the product maps $\Delta_{i}^{p},\Sigma_{i}^{p}$ also satisfy them componentwise, proving that $I_{\bullet}$ is a cosimplicial set.
2.4 The Diagonal Simplicial Module $X_{\bullet}(\vec{s};A)$
We define the contravariant functor $A^{(-)}:\mathbf{Set}\to\mathbf{Mod}_{A}$. This functor maps a set $S$ to the $A$-module of functions $S\to A$, and a map of sets $f:S_{1}\to S_{2}$ to the module homomorphism $A^{f}:A^{S_{2}}\to A^{S_{1}}$ defined by pre-composition ($T\mapsto T\circ f$).
Definition 2.2 (Diagonal simplicial tensor module).
The diagonal simplicial $A$-tensor module $X_{\bullet}(\vec{s};A):\Delta^{\mathrm{op}}\to\mathbf{Mod}_{A}$ is the composition $A^{(-)}\circ I_{\bullet}$.
On objects, $X_{p}(\vec{s};A):=A^{I_{p}}$. Since the index sets $I_{p}$ are finite, $X_{p}$ is a free $A$-module.
The face and degeneracy maps are induced by the corresponding maps in $I_{\bullet}$ via pre-composition:
$d_{i}^{p}(T)$ $:=T\circ\Delta_{i}^{p},$
$s_{i}^{p}(T)$ $:=T\circ\Sigma_{i}^{p}.$
Remark 2.3 (Matrices in the constant shape case).
If $k=2$ and $\vec{s}=(N,N)$ is constant, then n=N1n=N-1 and
$M_{a}(p)=n_{a}-1-n+p=p\qquad(a=1,2).$
Hence
$I_{p}=[p]^{2}\qquad\text{and}\qquad X_{p}(\vec{s};A)=A^{[p]^{2}}\cong A^{(p+1)\times(p+1)}.$
In particular,
$X_{N-1}(\vec{s};A)\cong A^{[N-1]^{2}}\cong A^{N\times N},$
so an $N\times N$ matrix may be viewed as a simplex in degree N1N-1 of $X_{\bullet}(\vec{s};A)$. This identification will be used when we regard adjacency matrices of graphs on $N$ vertices as elements of $X_{N-1}(\vec{s};A)$.
2.5 Axis Symmetries and Permutations of Tensor Factors
For a shape $\vec{s}=(n_{1},\dots,n_{k})$, define its stabilizer
$\operatorname{Stab}(\vec{s}):=\{\sigma\in S_{k}:n_{\sigma(a)}=n_{a}\ \text{for all }a\}.$
Each $\sigma\in\operatorname{Stab}(\vec{s})$ acts (on the left) on the index sets $I_{p}$ by permuting coordinates:
$\sigma\cdot(m_{1},\dots,m_{k}):=(m_{\sigma^{-1}(1)},\dots,m_{\sigma^{-1}(k)}).$
This induces a (left) action on the diagonal simplicial module $X_{\bullet}(\vec{s};A)$ by precomposition:
$(\sigma\cdot T)(m):=T(\sigma^{-1}\cdot m),\qquad T\in X_{p}(\vec{s};A),\ m\in I_{p}.$
Lemma 2.4 (Equivariance under $\operatorname{Stab}(\vec{s})$).
The action of $\operatorname{Stab}(\vec{s})$ on $I_{\bullet}$ commutes with the cosimplicial structure maps. Consequently, for every $\sigma\in\operatorname{Stab}(\vec{s})$ and all $p,i$ we have
$d_{i}^{p}(\sigma\cdot T)=\sigma\cdot d_{i}^{p}(T),\qquad s_{i}^{p}(\sigma\cdot T)=\sigma\cdot s_{i}^{p}(T).$
The boundary operators $\partial_{p}$ also commute with the $\operatorname{Stab}(\vec{s})$-action.
Proof.
Let $\sigma\in\operatorname{Stab}(\vec{s})$. By definition, $n_{a}=n_{\sigma(a)}$, which implies $n_{a}=n_{\sigma^{-1}(a)}$. This ensures $M_{a}(p)=M_{\sigma^{-1}(a)}(p)$ for all $a,p$.
We first verify the commutation on the index sets. Let $m\in I_{p-1}$. We show $\sigma\cdot\Delta_{i}^{p}(m)=\Delta_{i}^{p}(\sigma\cdot m)$. The $a$-th component of the left side is
$(\sigma\cdot\Delta_{i}^{p}(m))_{a}=(\Delta_{i}^{p}(m))_{\sigma^{-1}(a)}=\delta_{i}^{M_{\sigma^{-1}(a)}(p)}(m_{\sigma^{-1}(a)}).$
The $a$-th component of the right side is
$(\Delta_{i}^{p}(\sigma\cdot m))_{a}=\delta_{i}^{M_{a}(p)}((\sigma\cdot m)_{a})=\delta_{i}^{M_{a}(p)}(m_{\sigma^{-1}(a)}).$
Since $M_{a}(p)=M_{\sigma^{-1}(a)}(p)$, the components are equal. Thus $\sigma\circ\Delta_{i}^{p}=\Delta_{i}^{p}\circ\sigma$. The argument for $\Sigma_{i}^{p}$ is analogous.
Now we verify the equivariance on $X_{p}$. Let $T\in X_{p}$.
$d_{i}^{p}(\sigma\cdot T)$ $=(\sigma\cdot T)\circ\Delta_{i}^{p}$
$=(T\circ\sigma^{-1})\circ\Delta_{i}^{p}$
$=T\circ(\sigma^{-1}\circ\Delta_{i}^{p})$
$=T\circ(\Delta_{i}^{p}\circ\sigma^{-1})\quad(\text{by commutativity on }I_{\bullet})$
$=(T\circ\Delta_{i}^{p})\circ\sigma^{-1}$
$=\sigma\cdot(T\circ\Delta_{i}^{p})=\sigma\cdot d_{i}^{p}(T).$
The proof for $s_{i}^{p}$ is analogous. The statement for $\partial_{p}$ follows by linearity. ∎
Remark 2.5 (Constant shape and symmetry subcomplexes).
Suppose $\vec{s}=(N,\dots,N)$ is constant. Then $\operatorname{Stab}(\vec{s})\cong S_{k}$ acts on $X_{\bullet}(\vec{s};A)$ by permuting tensor axes. By Lemma 2.4, the symmetric and alternating subspaces defined below are preserved by all structure maps and hence form simplicial $A$-submodules.
In particular, the symmetric and alternating subspaces
$X_{\bullet}^{\mathrm{Sym}}:=\{T:\sigma\cdot T=T\ \forall\sigma\in S_{k}\},\qquad X_{\bullet}^{\mathrm{Alt}}:=\{T:\sigma\cdot T=\operatorname{sgn}(\sigma)\,T\}$
form simplicial $A$-submodules. (If $k!$ is invertible in $A$, these are direct summands.) In A we show that the global contracting homotopy $H$ is $\operatorname{Stab}(\vec{s})$-equivariant, so it restricts to chain contractions on $X_{\bullet}^{\mathrm{Sym}}$ and $X_{\bullet}^{\mathrm{Alt}}$. Hence these symmetry subcomplexes are also contractible.
Remark 2.6 (Matrix and Hermitian transpose).
For $k=2$ and $\vec{s}=(N,N)$, the nontrivial element $\tau\in S_{2}$ acts by swapping the tensor axes, corresponding to the transpose: $(\tau\cdot T)(m_{1},m_{2})=T(m_{2},m_{1})$. By Lemma 2.4, all structure maps and the boundary operators $\partial_{p}$ commute with this action.
Over $K=\mathbb{C}$, viewing $X_{\bullet}(\vec{s};K)$ as a simplicial $\mathbb{R}$-module, complex conjugation commutes with all structure maps. The Hermitian conjugate action (transpose combined with conjugation) also commutes with $d_{i}^{p},s_{i}^{p}$ and $\partial_{p}$ (as $\mathbb{R}$-linear operators). The real subspaces of symmetric, skew-symmetric, Hermitian, and skew-Hermitian tensors form simplicial $\mathbb{R}$-submodules, and the global contracting homotopy $H$ restricts to each of them.
2.6 Base Change and Realization Maps
Proposition 2.7 (Base change for $X_{\bullet}(\vec{s};-)$).
For any ring homomorphism $\varphi:A\to B$ and shape $\vec{s}$, there is a canonical isomorphism of simplicial $B$-modules
$X_{\bullet}(\vec{s};A)\otimes_{A}B\;\xrightarrow[\cong]{\theta_{\bullet}}\;X_{\bullet}(\vec{s};B).$
Proof.
Each $X_{p}(\vec{s};A)=A^{I_{p}}$ is a finite free $A$-module with basis $\{E_{m}\}_{m\in I_{p}}$ (indicator functions). Define
$\theta_{p}:X_{p}(\vec{s};A)\otimes_{A}B\longrightarrow X_{p}(\vec{s};B),\qquad\theta_{p}(E_{m}\otimes b)=b\cdot E_{m}.$
Since $I_{p}$ is finite, $(A^{I_{p}})\otimes_{A}B\cong B^{I_{p}}$, so each $\theta_{p}$ is a $B$-module isomorphism.
Simplicial compatibility follows because the face and degeneracy maps are given by precomposition with the same maps $\Delta_{i}^{p},\Sigma_{i}^{p}$ on $I_{\bullet}$ in both $X_{\bullet}(\vec{s};A)$ and $X_{\bullet}(\vec{s};B)$, so each $\theta_{p}$ commutes with all $d_{i}^{p}$ and $s_{i}^{p}$. ∎
Corollary 2.8 (Naturality and flat base change).
Let $T\in X_{n}(\vec{s};A)$, and let $f_{T}:A[\Delta^{n}]\to X_{\bullet}(\vec{s};A)$ be the corresponding realization map (defined in Section 8). Let $\varphi:A\to B$ be a ring homomorphism, and let $T_{B}:=\theta_{n}(T\otimes 1)\in X_{n}(\vec{s};B)$. Let $\varphi_{p}:A[\Delta^{n}]_{p}\otimes_{A}B\xrightarrow{\cong}B[\Delta^{n}]_{p}$ denote the canonical isomorphism.
  1. 1.
    (Naturality) The following diagram commutes for all $p$:
    $A[\Delta^{n}]_{p}\otimes_{A}B$$X_{p}(\vec{s};A)\otimes_{A}B$$B[\Delta^{n}]_{p}$$X_{p}(\vec{s};B)$$\scriptstyle{f_{T,p}\otimes 1_{B}}$$\scriptstyle{\varphi_{p}}\scriptstyle{\cong}$$\scriptstyle{\theta_{p}}\scriptstyle{\cong}$$\scriptstyle{f_{T_{B},p}}$
  2. 2.
    (Kernel preservation) If $B$ is a flat $A$-module, the kernel sequence is preserved:
    $(\ker f_{T,p})\otimes_{A}B\;\cong\;\ker(f_{T_{B},p})\qquad(\forall p).$
Proof.
(1) Let $\alpha\in\Delta([p],[n])$ be a basis element of $A[\Delta^{n}]_{p}$. Both paths in the diagram map the element $\alpha\otimes b$ to $b\cdot f_{T_{B},p}(\alpha)$, establishing commutativity.
(2) Consider the exact sequence $0\to\ker f_{T,p}\to A[\Delta^{n}]_{p}\xrightarrow{f_{T,p}}X_{p}(\vec{s};A)$. If $B$ is a flat $A$-module, the functor $(-)\otimes_{A}B$ is exact. Applying it yields the exact sequence
$0\to(\ker f_{T,p})\otimes_{A}B\to A[\Delta^{n}]_{p}\otimes_{A}B\xrightarrow{f_{T,p}\otimes 1_{B}}X_{p}(\vec{s};A)\otimes_{A}B.$
This implies $(\ker f_{T,p})\otimes_{A}B\cong\ker(f_{T,p}\otimes 1_{B})$. By the commutativity of the diagram in (1) and the fact that $\varphi_{p}$ and $\theta_{p}$ are isomorphisms, $\ker(f_{T,p}\otimes 1_{B})$ is isomorphic to $\ker(f_{T_{B},p})$. ∎
3 Horns, Kernels, and Missing Indices
We investigate the structure of face kernels and the uniqueness of fillers for horns in $X_{\bullet}(\vec{s};A)$.
3.1 Horns and the Horn Kernel
Let $p\geq 1$ and $0\leq j\leq p$. We denote $F_{j}=[p]\setminus\{j\}$.
Definition 3.1 ($(p,j)$-Horn and Filler).
A $(p,j)$-horn in $X_{\bullet}$ is a tuple $H=(h_{i})_{i\in F_{j}}$ of elements $h_{i}\in X_{p-1}$ satisfying the horn compatibility condition:
$d_{i}(h_{\ell})=d_{\ell-1}(h_{i})\quad\text{for all }i<\ell\text{ with }i,\ell\in F_{j}.$
A tensor $T\in X_{p}$ is a filler of $H$ if $d_{i}(T)=h_{i}$ for all $i\in F_{j}$.
In the category of simplicial modules, every horn has at least one filler. The uniqueness of fillers is measured by the horn kernel.
Definition 3.2 (Horn Kernel).
The $(p,j)$-horn kernel $R_{p,j}$ is the submodule of $X_{p}$ defined by
$R_{p,j}:=\bigcap_{i\in F_{j}}\ker(d_{i}^{p}).$
If $T$ and $T^{\prime}$ are two fillers for the same horn $H$, their difference $T-T^{\prime}\in R_{p,j}$. The set of fillers $\mathcal{F}(H)$ is an affine space (torsor) modeled on $R_{p,j}$. Fillers are unique if and only if $R_{p,j}=\{0\}$.
3.2 Structure of the Face Kernels
We characterize the kernels of the face maps using the standard basis $\{E_{m}\}_{m\in I_{p}}$ of $X_{p}$.
Lemma 3.3 (Action on basis elements).
Let $m\in I_{p}$. The action of the face map $d_{i}^{p}$ on $E_{m}$ is
[…]
Proof.
Evaluate at $m^{\prime\prime}\in I_{p-1}$: $(d_{i}^{p}(E_{m}))(m^{\prime\prime})=E_{m}(\Delta_{i}^{p}(m^{\prime\prime}))=\delta_{m,\Delta_{i}^{p}(m^{\prime\prime})}$. If $m\notin\operatorname{im}(\Delta_{i}^{p})$, this is zero. If $m\in\operatorname{im}(\Delta_{i}^{p})$, since $\Delta_{i}^{p}$ is injective, there is a unique $m^{\prime}$ such that $\Delta_{i}^{p}(m^{\prime})=m$. Then $d_{i}^{p}(E_{m})=E_{m^{\prime}}$. ∎
Lemma 3.4 (Injectivity Lemma).
If $d_{i}^{p}(E_{m_{1}})=d_{i}^{p}(E_{m_{2}})\neq 0$, then $m_{1}=m_{2}$.
Proof.
By Lemma 3.3, the common value is some $E_{m^{*}}\in X_{p-1}$. Then $m_{1}=\Delta_{i}^{p}(m^{*})$ and $m_{2}=\Delta_{i}^{p}(m^{*})$. ∎
Theorem 3.5 (Basis of the Face Kernel).
The kernel of $d_{i}^{p}:X_{p}\to X_{p-1}$ is the free $A$-submodule
$\ker(d_{i}^{p})=\operatorname{span}_{A}\{E_{m}\mid m\in I_{p}\setminus\operatorname{im}(\Delta_{i}^{p})\}.$
Proof.
Let $T=\sum_{m\in I_{p}}a_{m}E_{m}\in X_{p}$. Applying $d_{i}^{p}$ gives
$d_{i}^{p}(T)=\sum_{m\in\operatorname{im}(\Delta_{i}^{p})}a_{m}d_{i}^{p}(E_{m}),$
since $d_{i}^{p}(E_{m})=0$ if $m\notin\operatorname{im}(\Delta_{i}^{p})$ (Lemma 3.3). By the Injectivity Lemma 3.4, the set $\{d_{i}^{p}(E_{m})\mid m\in\operatorname{im}(\Delta_{i}^{p})\}$ consists of distinct standard basis elements in $X_{p-1}$, hence is AA-linearly independent.
Thus $d_{i}^{p}(T)=0$ if and only if $a_{m}=0$ for all $m\in\operatorname{im}(\Delta_{i}^{p})$. ∎
3.3 Missing Indices and the Support Characterization
Definition 3.6 (Missing Index).
An index $m\in I_{p}$ is missing from the $(p,j)$-horn if
$\forall i\in F_{j}:m\notin\operatorname{im}(\Delta_{i}^{p}).$
Let $M_{p,j}\subset I_{p}$ denote the set of missing indices.
Theorem 3.7 (Support Characterization / Horn Kernel Basis).
The horn kernel $R_{p,j}$ is a free $A$-module with basis $\{E_{m}\}_{m\in M_{p,j}}$. Consequently, if $T_{1}$ and $T_{2}$ are two fillers of the same horn, their difference is supported on the missing indices: $\operatorname{supp}(T_{1}-T_{2})\subseteq M_{p,j}$.
Proof.
By Theorem 3.5, $\ker(d_{i}^{p})$ is the coordinate subspace spanned by the basis $B_{i}=\{E_{m}\mid m\notin\operatorname{im}(\Delta_{i}^{p})\}$. The horn kernel $R_{p,j}$ is the intersection of these coordinate subspaces for $i\in F_{j}$. The intersection of coordinate subspaces is spanned by the intersection of their bases, $\bigcap_{i\in F_{j}}B_{i}$. An element $E_{m}$ lies in this intersection if and only if $m\notin\operatorname{im}(\Delta_{i}^{p})$ for all $i\in F_{j}$, which defines $M_{p,j}$. The consequence regarding the support of the difference follows because $T_{1}-T_{2}\in R_{p,j}$. ∎
3.4 Combinatorial Characterization
We provide a combinatorial interpretation of missing indices. We denote the image of a multi-index $m=(m_{1},\dots,m_{k})$ by $\operatorname{im}(m)=\{m_{1},\dots,m_{k}\}$.
Proposition 3.8 (Characterization of Missing Indices).
An index $m\in I_{p}$ is missing from the $(p,j)$-horn ($m\in M_{p,j}$) if and only if $\operatorname{im}(m)\supseteq F_{j}$.
Proof.
We first establish the equivalence: $m\in\operatorname{im}(\Delta_{i}^{p})\iff i\notin\operatorname{im}(m)$. Recall $\Delta_{i}^{p}=\prod_{a=1}^{k}\delta_{i}^{M_{a}(p)}$. The image of $\delta_{i}^{q}$ is $q]\setminus\{i\}$.
$m\in\operatorname{im}(\Delta_{i}^{p})$ $\iff\forall a:m_{a}\in\operatorname{im}(\delta_{i}^{M_{a}(p)})$
$\iff\forall a:m_{a}\neq i\quad(\text{since }i\in[p]\subseteq[M_{a}(p)])$
$\iff i\notin\operatorname{im}(m).$
Taking the contrapositive, $m\notin\operatorname{im}(\Delta_{i}^{p})\iff i\in\operatorname{im}(m)$.
By definition, $m\in M_{p,j}$ if and only if $\forall i\in F_{j}:m\notin\operatorname{im}(\Delta_{i}^{p})$. Therefore,
$m\in M_{p,j}\iff\forall i\in F_{j}:i\in\operatorname{im}(m)\iff F_{j}\subseteq\operatorname{im}(m).$
Corollary 3.9.
The horn kernel is non-trivial ($R_{p,j}\neq\{0\}$) if and only if $k\geq p$.
Proof.
By Theorem 3.7, $R_{p,j}\neq\{0\}$ iff $M_{p,j}\neq\emptyset$.
($\Rightarrow$) If $m\in M_{p,j}$, then $k\geq|\operatorname{im}(m)|\geq|F_{j}|=p$.
($\Leftarrow$) Suppose $k\geq p$. Let $F_{j}=\{f_{1},\dots,f_{p}\}$. Define $m=(f_{1},\dots,f_{p},0,\dots,0)\in\mathbb{N}^{k}$. Since $M_{a}(p)\geq p$, we have $p]^{k}\subseteq I_{p}$, so $m\in I_{p}$. Since $\operatorname{im}(m)\supseteq F_{j}$, $m\in M_{p,j}$. ∎
3.5 The $\ell$-free Subspace
Fix a degree $p\geq 0$ and an index $0\leq\ell\leq p$. The combinatorial characterization of $\operatorname{im}(\Delta_{\ell}^{p})$ yields a structural property of the face maps.
Definition 3.10 ($\ell$-free subspace).
Define the $\ell$-free subspace
$X_{p}^{(\ell\text{-free})}:=\operatorname{span}_{A}\{\,E_{m}:\ \ell\notin\operatorname{im}(m)\,\}\ \subseteq X_{p}.$
Corollary 3.11 (Isomorphism on the $\ell$-free subspace).
The face map $d_{\ell}^{p}$ induces a linear isomorphism
$d_{\ell}^{p}:\ X_{p}^{(\ell\text{-free})}\ \xrightarrow{\ \cong\ }\ X_{p-1}.$
Proof.
By the proof of Proposition 3.8, $m\in\operatorname{im}(\Delta_{\ell}^{p})\iff\ell\notin\operatorname{im}(m)$. Thus $X_{p}^{(\ell\text{-free})}$ is spanned by $\{E_{m}\mid m\in\operatorname{im}(\Delta_{\ell}^{p})\}$. By the Injectivity Lemma 3.4, $d_{\ell}^{p}$ is injective on this subspace. Since $\Delta_{\ell}^{p}:I_{p-1}\to\operatorname{im}(\Delta_{\ell}^{p})$ is a bijection, Lemma 3.3 shows that $d_{\ell}^{p}$ maps the basis of $X_{p}^{(\ell\text{-free})}$ bijectively onto the basis of $X_{p-1}$. ∎
3.6 Genericity
Let $T^{\prime}=\mu_{j}(\Phi_{j}(T))$ be the Moore filler of $T$ (see Section 4). The difference $T-T^{\prime}$ lies in $R_{p,j}$. By Theorem 3.7, $\operatorname{supp}(T-T^{\prime})\subseteq M_{p,j}$.
Definition 3.12 (Generic Tensor).
Assume $A$ is an integral domain. A tensor $T\in X_{p}$ is called $(p,j)$-generic if the difference $T-T^{\prime}$ has maximal support: $\operatorname{supp}(T-T^{\prime})=M_{p,j}$. (This requires $k\geq p$).
Proposition 3.13 (Generic Locus).
Assume $A$ is an integral domain and $k\geq p$. The set of $(p,j)$-generic tensors forms a Zariski-open subset $\mathcal{U}_{p,j}\subset X_{p}$. If moreover $A$ is infinite, then $\mathcal{U}_{p,j}$ is non-empty.
Proof.
Let $R(T)=T-T^{\prime}$. The map $T\mapsto T^{\prime}$ is $A$-linear (as $\mu_{j}$ and $\Phi_{j}$ are linear), hence $R(T)$ is $A$-linear. Write $R(T)=\sum_{m\in M_{p,j}}c_{m}(T)E_{m}$. Each coefficient $c_{m}(T)$ is a linear form in the entries of $T$.
We show that $c_{m}(T)$ is not identically zero. Consider $T=E_{m}$ for $m\in M_{p,j}$. Since $E_{m}\in R_{p,j}$, its horn is $\Phi_{j}(E_{m})=0$. The Moore filler of the zero horn is $T^{\prime}=\mu_{j}(0)=0$. Then $R(E_{m})=E_{m}$, so $c_{m}(E_{m})=1$.
The generic locus is defined by the non-vanishing of these forms:
$\mathcal{U}_{p,j}\ :=\ \{\,T\in X_{p}:\ c_{m}(T)\neq 0\text{ for all }m\in M_{p,j}\,\}.$
The complement is the union of the zero loci of the non-zero forms $c_{m}$, hence a finite union of proper linear subspaces. Therefore $\mathcal{U}_{p,j}$ is Zariski-open. If $A$ is infinite, a finite union of proper linear subspaces cannot equal $X_{p}$, so $\mathcal{U}_{p,j}$ is non-empty. ∎
4 Normalization and the Moore Filler
We review the normalization of simplicial modules. Let $X_{\bullet}$ be a simplicial $A$-module.
4.1 The Normalization Theorem
Definition 4.1 (Normalized and degenerate submodules).
The normalized submodule in degree $p$ is $N_{p}(X):=\bigcap_{i=1}^{p}\ker d_{i}\subseteq X_{p}$. The degenerate submodule in degree $p$ is $D_{p}(X):=\sum_{r=0}^{p-1}\operatorname{im}(s_{r})\subseteq X_{p}$.
By the normalization theorem, the homology of $X_{\bullet}$ is isomorphic to the homology of its normalized (Moore) complex $(N_{\bullet}(X),d_{0})$.
Theorem 4.2 (Normalization theorem / Dold-Kan Correspondence).
For every $p\geq 0$ there is a functorial direct sum decomposition
$X_{p}\cong N_{p}(X)\oplus D_{p}(X).$
There is a functorial projection $\pi_{p}:X_{p}\to N_{p}(X)$ with $\ker(\pi_{p})=D_{p}(X)$. Moreover, $X_{p}$ decomposes as a direct sum of images of $N_{q}(X)$ under iterated degeneracies (Eilenberg–Zilber decomposition):
$X_{p}\;=\;\bigoplus_{q=0}^{p}\ \bigoplus_{\begin{subarray}{c}0\leq i_{1}<\cdots<i_{p-q}\leq p-1\end{subarray}}s_{i_{p-q}}\cdots s_{i_{1}}\,N_{q}(X).$
Refs.: Dold [Dold1958]; Kan [Kan1958]; May [MaySimplicial]; Weibel [Weibel, Thm 8.3.8, Cor 8.4.2]; Goerss–Jardine [GoerssJardine, III.2].
Remark 4.3 (Eilenberg–Mac Lane idempotents).
The decomposition is established by the Eilenberg–Mac Lane idempotents. The projection onto $N_{p}(X)$ corresponding to the convention $(N_{\bullet},d_{0})$ is given explicitly by $\pi_{p}=(\operatorname{id}-s_{p-1}d_{p})\cdots(\operatorname{id}-s_{0}d_{1})$. These operators rely only on the simplicial identities and are thus functorial for any simplicial module. Refs.: Mac Lane [MacLaneHomology, Ch. VIII, §6]; Weibel [Weibel, Thm 8.3.8].
4.2 The Moore Horn Filler Algorithm
Definition 4.4 (Horn space).
Fix $p\geq 1$ and $j\in\{0,\dots,p\}$. Let $F_{j}=[p]\setminus\{j\}$. The space of compatible $(p,j)$-horns, $\operatorname{Horns}(p,j)$, is the submodule of $\prod_{i\in F_{j}}X_{p-1}$ defined by the compatibility conditions $d_{i}x_{\ell}=d_{\ell-1}x_{i}$ for $i<\ell$ in $F_{j}$.
The horn map is $\Phi_{j}:X_{p}\to\prod_{i\in F_{j}}X_{p-1}$, defined by $\Phi_{j}(T)=(d_{i}T)_{i\in F_{j}}$. Its image lies in $\operatorname{Horns}(p,j)$, and its kernel is the horn kernel $R_{p,j}(X)=\bigcap_{i\in F_{j}}\ker d_{i}$.
Proposition 4.5 (Exactness of the Horn Sequence).
The sequence
$0\longrightarrow R_{p,j}(X)\longrightarrow X_{p}\xrightarrow{\ \Phi_{j}\ }\operatorname{Horns}(p,j)\longrightarrow 0$
is exact. The surjectivity of $\Phi_{j}$ (the Kan condition in the abelian setting) is established by an explicit iterative construction (the Moore filler algorithm) which yields a filler in the degenerate submodule $D_{p}(X)$.
Proof.
Exactness at $R_{p,j}(X)$ and $X_{p}$ is by definition. The surjectivity of $\Phi_{j}$ and the construction of a degenerate filler is standard; see e.g., Weibel [Weibel, Lemma 8.2.6] or Duskin [Duskin1979, Lemma 3.1]. ∎
We detail the algorithm used in the proof of Proposition 4.5.
Definition 4.6 (Moore Filler Map).
The Moore filler map $\mu_{j}:\operatorname{Horns}(p,j)\to D_{p}(X)$ is defined iteratively for a horn $H=(x_{i})_{i\in F_{j}}$.
Phase I (Ascending indices $i<j$). Initialize $T^{(-1)}:=0$. For $i=0,\dots,j-1$:
$T^{(i)}:=T^{(i-1)}+s_{i}\bigl(x_{i}-d_{i}T^{(i-1)}\bigr).$
Phase II (Descending indices $i>j$). Initialize $U^{(p+1)}:=T^{(j-1)}$. For $i=p,\dots,j+1$:
$U^{(i)}:=U^{(i+1)}+s_{i-1}\bigl(x_{i}-d_{i}U^{(i+1)}\bigr).$
Output. $\mu_{j}(H):=U^{(j+1)}$.
By construction, $\mu_{j}(H)\in D_{p}(X)$. The map $\mu_{j}$ provides an explicit right inverse (section) to $\Phi_{j}$.
Corollary 4.7.
The Moore filler map $\mu_{j}$ induces a direct sum decomposition
$X_{p}=R_{p,j}(X)\oplus\operatorname{im}(\mu_{j}).$
Furthermore, $\operatorname{im}(\mu_{j})\subseteq D_{p}(X)$.
Proof.
Since $\mu_{j}$ is a section of $\Phi_{j}$ ($\Phi_{j}\circ\mu_{j}=\mathrm{id}$), the exact sequence in Proposition 4.5 splits, yielding the decomposition. The inclusion follows from Definition 4.6. ∎
5 Combinatorics and Classification
We apply the combinatorial characterization of the horn kernel (Section 3) to classify the DSTM $X_{\bullet}(\vec{s};A)$ and derive exact rank formulas.
5.1 Classification Theorem
We establish the main classification result based on the tensor order $k$ and the simplicial dimension $n=\min(\vec{s})-1$.
Definition 5.1 (Strict Algebraic $n$-Hypergroupoid).
A simplicial module $X_{\bullet}$ is a strict algebraic $n$-hypergroupoid if fillers are unique in dimensions strictly greater than $n$, and not unique in dimension $n$. That is:
  1. 1.
    $R_{p,j}(X)=0$ for all $p>n$ and all $j$.
  2. 2.
    $R_{n,j}(X)\neq 0$ for at least one $j$.
Remark 5.2 (Relation to Duskin–Glenn).
Condition (1) in Definition 5.1 is exactly the $n$-hypergroupoid condition of Duskin–Glenn (as in the nnLab definition): all horns in dimensions $p>n$ admit unique fillers. Condition (2) adds a non-uniqueness requirement in dimension $n$: there exists $j$ with $R_{n,j}(X)\neq 0$. Thus a strict algebraic $n$-hypergroupoid in our sense is an $n$-hypergroupoid in which horn fillers are non-unique in dimension $n$ and unique in all dimensions $p>n$.
Theorem 5.3 (Hypergroupoid Classification).
$X_{\bullet}(\vec{s};A)$ is a strict algebraic $n$-hypergroupoid if and only if $k=n$.
Proof.
We use the criterion established in Corollary 3.9: for each $p$ and $j$, $R_{p,j}\neq 0$ if and only if $k\geq p$.
Condition (1) requires $R_{p,j}=0$ for all $p>n$. This is equivalent to $k<p$ for all $p\geq n+1$. This holds if and only if $k<n+1$, i.e., $k\leq n$.
Condition (2) requires $R_{n,j}\neq 0$. This is equivalent to $k\geq n$.
$X_{\bullet}$ is a strict algebraic $n$-hypergroupoid if and only if both conditions hold: $(k\leq n)\land(k\geq n)$, which is equivalent to $k=n$. ∎
5.2 Rank Formulas via Inclusion-Exclusion
We analyze the ranks of the components of the DSTM. Since $X_{p}(\vec{s};A)$ is a free $A$-module and the normalization projections are defined over $\mathbb{Z}$ (Remark 4.3), the ranks of the submodules $R_{p,j},N_{p},D_{p}$ are independent of the ring $A$.
Let $N_{a}=n_{a}-1$. Recall $n=\min_{a}N_{a}$.
Proposition 5.4 (Rank of $R_{p,j}$).
Let $F_{j}=[p]\setminus\{j\}$. The rank of $R_{p,j}(X)$ is
$\operatorname{rank}R_{p,j}=\sum_{t=0}^{p}(-1)^{t}\binom{p}{t}\prod_{a=1}^{k}(M_{a}(p)+1-t).$
Proof.
By Theorem 3.7 and Proposition 3.8, the rank is the count of $m\in I_{p}$ such that $\operatorname{im}(m)\supseteq F_{j}$. We use inclusion–exclusion on $I_{p}=\prod[M_{a}(p)]$.
For a subset $S\subseteq F_{j}$ with $|S|=t$, we count the indices $m\in I_{p}$ such that $\operatorname{im}(m)\cap S=\emptyset$. Since $M_{a}(p)\geq p$, $S\subseteq[M_{a}(p)]$. The number of choices for the $a$-th component is $|[M_{a}(p)]\setminus S|=M_{a}(p)+1-t$. The total count of such indices is $\prod_{a=1}^{k}(M_{a}(p)+1-t)$.
Since $|F_{j}|=p$, the inclusion–exclusion principle yields the formula for the count of indices covering $F_{j}$. ∎
We analyze the normalized complex $(N_{\bullet},d_{0})$. The cycles are $Z_{p}(N_{\bullet})=\ker(d_{0}:N_{p}\to N_{p-1})=\bigcap_{i=0}^{p}\ker d_{i}$.
Corollary 5.5 (Rank of $Z_{p}(N_{\bullet})$).
The rank of the normalized cycles $Z_{p}(N_{\bullet})$ is
$\operatorname{rank}Z_{p}(N_{\bullet})=\sum_{t=0}^{p+1}(-1)^{t}\binom{p+1}{t}\prod_{a=1}^{k}(M_{a}(p)+1-t).$
Proof.
$Z_{p}(N_{\bullet})$ is spanned by basis elements $E_{m}$ such that $\operatorname{im}(m)\supseteq[p]$. The formula follows by applying inclusion–exclusion with the covered set $[p]$ (size $p+1$). ∎
5.3 Constant Shape and Stirling Numbers
If the shape is constant, then $n_{a}=n+1$ for all $a$, where $n$ is the simplicial dimension. In this case $N_{a}=n$ and hence $M_{a}(p)=p$ for all $a$.
[…]
Theorem 5.6 (Ranks for Constant Shape).
If $M_{a}(p)=p$ for all $a$ (i.e., $I_{p}=[p]^{k}$), the ranks of the components of the Moore complex in degree $p$ are:
[…]
Proof.
We interpret these ranks in terms of finite differences of the polynomial $P(x)=x^{k}$.
The rank of $Z_{p}(N_{\bullet})$ specializes to
$\sum_{t=0}^{p+1}(-1)^{t}\binom{p+1}{t}(p+1-t)^{k}=\Delta^{p+1}(x^{k})\big|_{x=0}.$
[…]
The rank of $N_{p}(X)=R_{p,0}$ specializes to
$\sum_{t=0}^{p}(-1)^{t}\binom{p}{t}(p+1-t)^{k}=\Delta^{p}(x^{k})\big|_{x=1}.$
[…]
5.4 Boundaries and Contractibility
We analyze the boundaries $B_{p}(N_{\bullet})=\operatorname{im}(d_{0}:N_{p+1}\to N_{p})$.
Proposition 5.7 (Non-triviality of Boundaries).
$B_{p}(N_{\bullet})\neq 0$ if and only if $k\geq p+1$.
Proof.
If $k<p+1$, then $N_{p+1}(X)=0$ by Corollary 3.9 (since $N_{p+1}=R_{p+1,0}$), so $B_{p}(N_{\bullet})=0$.
If $k\geq p+1$, we construct a non-zero boundary. We seek $T\in N_{p+1}(X)$ such that $d_{0}(T)\neq 0$. We require an index $m\in I_{p+1}$ such that $E_{m}\in N_{p+1}(X)$ and $d_{0}(E_{m})\neq 0$. $E_{m}\in N_{p+1}(X)$ requires $\operatorname{im}(m)\supseteq\{1,\dots,p+1\}$. $d_{0}(E_{m})\neq 0$ requires $m\in\operatorname{im}(\Delta_{0}^{p+1})$, which is equivalent to $0\notin\operatorname{im}(m)$. Thus we seek $m$ such that $\operatorname{im}(m)=\{1,\dots,p+1\}$. Since $k\geq p+1$, such an $m$ exists. Since $M_{a}(p+1)\geq p+1$, $m\in I_{p+1}$. Therefore $B_{p}(N_{\bullet})\neq 0$. ∎
The DSTM $X_{\bullet}$ is contractible (Theorem A.4). Consequently, $H_{*}(X_{\bullet})=0$, implying $Z_{p}(N_{\bullet})=B_{p}(N_{\bullet})$ for all $p$. The differential $d_{0}:N_{p+1}\to N_{p}$ maps surjectively onto $Z_{p}(N_{\bullet})$ with kernel $Z_{p+1}(N_{\bullet})$.
Proposition 5.8 (Rank Consistency Check).
The rank formulas derived satisfy the consistency condition required by contractibility (Rank-Nullity):
$\operatorname{Rank}N_{p+1}(X)=\operatorname{Rank}Z_{p+1}(N_{\bullet})+\operatorname{Rank}Z_{p}(N_{\bullet}).$
Proof.
We verify the identity using the formulas from Theorem 5.6 in the constant shape case.
[…]
This matches the formula for $\operatorname{Rank}N_{p+1}(X)$ derived in Theorem 5.6 (by replacing $p$ with $p+1$). The consistency for the general shape case follows because the complex is defined and contractible over $\mathbb{Z}$. ∎
6 The Horn Non-Degeneracy Lemma and Decomposition
We establish that the horn kernel and the degenerate submodule intersect trivially. This ensures the Horn decomposition is compatible with the standard decomposition of simplicial modules.
Lemma 6.1 (Horn Non-Degeneracy).
For any simplicial module $X_{\bullet}$, any degree $p\geq 1$, and any $j\in[p]$,
$R_{p,j}(X)\cap D_{p}(X)=\{0\}.$
Proof.
We filter the degenerate submodule $D_{p}(X)$ by the maximum index of the degeneracy operators. For $0\leq r\leq p-1$, set
$H_{r}:=\sum_{i=0}^{r}\operatorname{im}(s_{i})\subseteq X_{p},\quad H^{\prime}_{r}:=\sum_{i=0}^{r}\operatorname{im}(s_{i})\subseteq X_{p-1},$
and set $H_{-1}=H^{\prime}_{-1}=\{0\}$. Note that $D_{p}(X)=H_{p-1}$.
Let $x\in R_{p,j}(X)\cap D_{p}(X)$ and assume $x\neq 0$. Choose $m\geq 0$ minimal such that $x\in H_{m}$. Then we can write
$x=s_{m}(z)+y,\qquad z\in X_{p-1},\ \ y\in H_{m-1}.$
If we can show $z\in H^{\prime}_{m-1}$, then $z=\sum_{i=0}^{m-1}s_{i}(u_{i})$. Using the identity $s_{m}s_{i}=s_{i}s_{m-1}$ for $i<m$, we obtain
$s_{m}(z)=\sum_{i=0}^{m-1}s_{m}s_{i}(u_{i})=\sum_{i=0}^{m-1}s_{i}s_{m-1}(u_{i})\in H_{m-1}.$
This implies $x\in H_{m-1}$, contradicting the minimality of $m$. Thus, it suffices to prove $z\in H^{\prime}_{m-1}$.
We split the proof into cases based on the index $j$.
Case 1: $j\neq m+1$. Since $m+1\leq p$ and $m+1\neq j$, we have $d_{m+1}(x)=0$.
$0=d_{m+1}(x)=d_{m+1}s_{m}(z)+d_{m+1}(y)=z+d_{m+1}(y),$
where we used $d_{m+1}s_{m}=\mathrm{id}$. Since $y\in H_{m-1}$, we write $y=\sum_{i=0}^{m-1}s_{i}(w_{i})$. For $i<m$, the identity $d_{m+1}s_{i}=s_{i}d_{m}$ holds, so
$d_{m+1}(y)=\sum_{i=0}^{m-1}s_{i}d_{m}(w_{i})\in H^{\prime}_{m-1}.$
Hence $z=-d_{m+1}(y)\in H^{\prime}_{m-1}$, as required.
Case 2: $j=m+1$. In this case, the face maps $d_{0},\dots,d_{m}$ all annihilate $x$ (since $j=m+1$). Since $x\in H_{m}$, we may write $x$ as a sum $x=\sum_{i=0}^{m}s_{i}(w_{i})$ for some coefficients $w_{i}\in X_{p-1}$. (Note that $z$ from the setup corresponds to $w_{m}$).
We prove by induction on $k$ (for $0\leq k\leq m$) that $w_{k}\in H^{\prime}_{m-1}$.
Base step ($k=0$): Consider $d_{0}(x)=0$.
$0=\sum_{i=0}^{m}d_{0}s_{i}(w_{i})=d_{0}s_{0}(w_{0})+\sum_{i=1}^{m}d_{0}s_{i}(w_{i}).$
Using $d_{0}s_{0}=\mathrm{id}$ and $d_{0}s_{i}=s_{i-1}d_{0}$ for $i\geq 1$:
$0=w_{0}+\sum_{i=1}^{m}s_{i-1}d_{0}(w_{i}).$
The sum consists of terms in $\operatorname{im}(s)$, so $w_{0}\in H^{\prime}_{m-1}$.
Inductive step: Assume $w_{0},\dots,w_{k-1}\in H^{\prime}_{m-1}$ for some $1\leq k\leq m$. Consider $d_{k}(x)=0$.
$0=\sum_{i=0}^{m}d_{k}s_{i}(w_{i}).$
We split the sum into four parts based on the index $i$:
  1. 1.
    i<k1i<k-1: $d_{k}s_{i}(w_{i})=s_{i}d_{k-1}(w_{i})\in H^{\prime}_{m-1}$.
  2. 2.
    i=k1i=k-1: $d_{k}s_{k-1}(w_{k-1})=w_{k-1}$ (using $d_{k}s_{k-1}=\mathrm{id}$). By the inductive hypothesis, $w_{k-1}\in H^{\prime}_{m-1}$.
  3. 3.
    $i=k$: $d_{k}s_{k}(w_{k})=w_{k}$ (using $d_{k}s_{k}=\mathrm{id}$).
  4. 4.
    $i>k$: $d_{k}s_{i}(w_{i})=s_{i-1}d_{k}(w_{i})\in H^{\prime}_{m-1}$.
Substituting these into the equation yields
[…]
Solving for $w_{k}$, we see that $w_{k}$ is a sum of elements in $H^{\prime}_{m-1}$. Thus $w_{k}\in H^{\prime}_{m-1}$.
By induction, $w_{m}\in H^{\prime}_{m-1}$. Since $z=w_{m}$, we have $z\in H^{\prime}_{m-1}$.
In all cases, we conclude $z\in H^{\prime}_{m-1}$, which contradicts the minimality of $m$ unless $x=0$. ∎
Theorem 6.2 (Horn decomposition).
For any simplicial module $X_{\bullet}$, any $p\geq 1$, and any $j\in[p]$, the Moore filler map $\mu_{j}$ (Definition 4.6) is a section of the horn map $\Phi_{j}$, and
$X_{p}=R_{p,j}(X)\oplus\operatorname{im}(\mu_{j}),\qquad\operatorname{im}(\mu_{j})\subseteq D_{p}(X).$
In particular, the horn kernel $R_{p,j}(X)$ is a direct summand of $X_{p}$, complementary to a degenerate summand.
Proof.
By construction (Definition 4.6) and Proposition 4.5, $\mu_{j}$ is a section of $\Phi_{j}$, i.e. $\Phi_{j}\circ\mu_{j}=\mathrm{id}_{\operatorname{Horns}(p,j)}$. Thus the short exact sequence in Proposition 4.5 splits, and we obtain
$X_{p}=R_{p,j}(X)\oplus\operatorname{im}(\mu_{j}).$
Each step in the definition of $\mu_{j}$ adds a degeneracy, so $\operatorname{im}(\mu_{j})\subseteq D_{p}(X)$. Lemma 6.1 ensures that the intersection of these summands is trivial ($R_{p,j}(X)\cap\operatorname{im}(\mu_{j})=\{0\}$), confirming that the decomposition is compatible with the standard degenerate filtration. ∎
Remark 6.3 (Normalization Conventions).
Theorem 6.2 generalizes the classical Normalization Theorem (Theorem 4.2). Standard conventions for the Moore complex (or normalized chain complex) vary:
  1. 1.
    $N_{p}^{(0)}(X)=\bigcap_{i=0}^{p-1}\ker d_{i}=R_{p,p}(X)$.
  2. 2.
    $N_{p}^{(1)}(X)=\bigcap_{i=1}^{p}\ker d_{i}=R_{p,0}(X)$.
In this paper we adopt the second convention $N_{p}(X)=N_{p}^{(1)}(X)=R_{p,0}(X)$. Our horn decomposition encompasses both as the cases $j=p$ and $j=0$ respectively.
7 Homology Dichotomy and Classification
7.1 The Horn Complex
In this subsection we assume $n\geq 2$; for $n<2$ the horn complex in low degrees degenerates and the corresponding homology statements are easily checked directly. Fix $j\in[n]$. Define the linear maps
$\Phi_{j}:X_{n}\longrightarrow\bigoplus_{i\neq j}X_{n-1},\qquad\Phi_{j}(T)=(d_{i}T)_{i\neq j},$
$\Psi:\bigoplus_{i\neq j}X_{n-1}\longrightarrow\bigoplus_{\begin{subarray}{c}i<m\\ i,m\neq j\end{subarray}}X_{n-2},\qquad\Psi((x_{i}))=(d_{i}x_{m}-d_{m-1}x_{i})_{i<m}.$
Proposition 7.1.
The sequence
$C^{\mathrm{horn}}_{j}:\qquad 0\longrightarrow X_{n}\xrightarrow{\ \partial_{2}=\Phi_{j}\ }\bigoplus_{i\neq j}X_{n-1}\xrightarrow{\ \partial_{1}=\Psi\ }\bigoplus_{i<m,\,i,m\neq j}X_{n-2}\longrightarrow 0$
is a chain complex (indexed homologically with $X_{n}$ in degree $2$). Moreover,
$H_{2}\bigl(C^{\mathrm{horn}}_{j}\bigr)\ \cong\ R_{n,j},\qquad H_{1}\bigl(C^{\mathrm{horn}}_{j}\bigr)=0.$
Proof.
$\partial_{1}\circ\partial_{2}=0$ by the simplicial identities $d_{i}d_{m}=d_{m-1}d_{i}$ for $i<m$. Since the complex starts in degree $2$, $H_{2}=\ker\partial_{2}=R_{n,j}$. The kernel of $\partial_{1}$ is the space of compatible horns $\operatorname{Horns}(n,j)$; by the existence of fillers (Proposition 4.5), $\operatorname{im}\partial_{2}=\operatorname{Horns}(n,j)$. Thus $H_{1}=\ker\partial_{1}/\operatorname{im}\partial_{2}=0$. ∎
Theorem 7.2 (Short exact sequence in the horn kernel).
The face map $d_{j}:R_{n,j}\to X_{n-1}$ induces a natural short exact sequence
$0\longrightarrow Z_{n}\longrightarrow R_{n,j}\xrightarrow{\,d_{j}\,}d_{j}(R_{n,j})\longrightarrow 0.$
Proof.
The kernel of $d_{j}|_{R_{n,j}}$ is $R_{n,j}\cap\ker d_{j}=\bigcap_{i\neq j}\ker d_{i}\cap\ker d_{j}=Z_{n}$. ∎
7.2 The Classification Dichotomy
We use the ranks established in Section 5.
Corollary 7.3 (Filler dichotomy).
The ranks of $R_{n,j}$ and $Z_{n}$ (Proposition 5.4, Corollary 5.5) determine the following classification based on the tensor order $k$:
  1. 1.
    If $k<n$, then $R_{n,j}=0$ (fillers are unique in dimension $n$).
  2. 2.
    If $k=n$, then $R_{n,j}\neq 0$ and $Z_{n}=0$. Hence $R_{n,j}\cong d_{j}(R_{n,j})$.
  3. 3.
    If $k\geq n+1$, then $Z_{n}\neq 0$ and $Z_{n}$ injects canonically into $R_{n,j}$.
Corollary 7.4 (Constant shape ranks).
For the constant shape $\vec{s}=(n+1,\ldots,n+1)$, the ranks are given by Theorem 5.6:
[…]
In particular:
  • 1.
    $Z_{n}\neq 0$ if and only if $k\geq n+1$.
  • 2.
    $R_{n,j}\neq 0$ if and only if $k\geq n$.
Remark 7.5 (Moore filler vs. $T$).
For the horn $\Lambda^{n}_{j}(T)$, Moore’s filler $T^{\prime}$ satisfies $T^{\prime}=T$ if and only if $T\in D_{n}$: the “only if” direction holds because $T^{\prime}$ is always degenerate (Definition 4.6), and the “if” direction follows from Lemma 6.1, since in that case $T-T^{\prime}$ lies in $R_{n,j}\cap D_{n}=\{0\}$. If $k<n$, then $R_{n,j}=0$, so $T^{\prime}=T$ always holds (and $D_{n}=X_{n}$). If $k\geq n$ and $T\notin D_{n}$, then $T-T^{\prime}\in R_{n,j}\setminus\{0\}$ and is supported on the missing indices (Theorem 3.7).
7.3 Interpretation as Algebraic $n$-Hypergroupoids
We adapt the definition of Duskin (1979) and Glenn (1982).
Definition 7.6 (Algebraic $n$-hypergroupoid).
A simplicial module $X_{\bullet}$ is an algebraic $n$-hypergroupoid if horn fillers are unique in dimensions $p>n$. It is strict if it is not an (n1)(n-1)-hypergroupoid (i.e., fillers in dimension $n$ are not unique).
Proposition 7.7.
A simplicial module $X_{\bullet}$ is an algebraic $n$-hypergroupoid if and only if $R_{p,j}(X)=0$ for all $p>n$ and all $j$.
Proof.
Simplicial modules are Kan complexes (fillers always exist, e.g., Proposition 4.5). As observed in Section 3, the set of fillers of a given horn is an affine torsor modeled on $R_{p,j}$, hence a singleton if and only if $R_{p,j}=0$. ∎
Applying the criterion of Corollary 3.9 to the DSTM $X_{\bullet}(\vec{s};A)$ with simplicial dimension $n$ yields the hypergroupoid classification of Theorem 5.3: $X_{\bullet}(\vec{s};A)$ is a strict algebraic $n$-hypergroupoid if and only if the tensor order $k$ equals $n$.
7.4 The Subcomplex of Normalized Cycles
We analyze the subcomplex formed by the normalized cycles $Z_{r}(N_{\bullet})$. In the DSTM, $Z_{r}(N_{\bullet})$ is spanned by basis elements $E_{m}$ such that $\operatorname{im}(m)\supseteq[r]$.
Lemma 7.8.
The family $(Z_{r}(N_{\bullet}))_{r\geq 0}$ forms a chain subcomplex of $(X_{\bullet},\partial)$. Furthermore, the differential on this subcomplex is zero:
$\partial_{r}\bigl(Z_{r}(N_{\bullet})\bigr)=0,\qquad H_{r}\!\bigl(Z_{\bullet}(N_{\bullet})\bigr)\cong Z_{r}(N_{\bullet}).$
There is a short exact sequence of chain complexes
$0\longrightarrow Z_{\bullet}(N_{\bullet})\longrightarrow X_{\bullet}\longrightarrow X_{\bullet}/Z_{\bullet}(N_{\bullet})\longrightarrow 0.$ (1)
Proof.
By definition, $Z_{r}(N_{\bullet})=\bigcap_{i=0}^{r}\ker d_{i}$. If $x\in Z_{r}(N_{\bullet})$, then $d_{i}(x)=0$ for all $i$. Therefore the boundary $\partial_{r}(x)=\sum(-1)^{i}d_{i}(x)=0$. This confirms that $Z_{\bullet}(N_{\bullet})$ is a chain subcomplex with zero differential. The homology of a complex with zero differential is the complex itself. ∎
Corollary 7.9 (Quotient homology).
For the DSTM $X_{\bullet}(\vec{s};A)$, which is contractible by Theorem A.4, the short exact sequence of chain complexes (1) induces isomorphisms
$H_{r}\!\bigl(X_{\bullet}/Z_{\bullet}(N_{\bullet})\bigr)\ \cong\ H_{r-1}\!\bigl(Z_{\bullet}(N_{\bullet})\bigr)\cong Z_{r-1}(N_{\bullet})$
for all $r\geq 1$.
Proof.
By Theorem A.4, the DSTM $X_{\bullet}(\vec{s};A)$ is contractible, hence $H_{r}(X_{\bullet})=0$ for all $r\geq 0$. Taking homology of the short exact sequence (1) gives a long exact sequence
$\cdots\to H_{r}(Z_{\bullet}(N_{\bullet}))\to H_{r}(X_{\bullet})\to H_{r}(X_{\bullet}/Z_{\bullet}(N_{\bullet}))\to H_{r-1}(Z_{\bullet}(N_{\bullet}))\to\cdots.$
The differential on $Z_{\bullet}(N_{\bullet})$ is zero, so $H_{r}(Z_{\bullet}(N_{\bullet}))\cong Z_{r}(N_{\bullet})$ for all $r$. Using $H_{r}(X_{\bullet})=0$ and exactness, we obtain
$H_{r}\!\bigl(X_{\bullet}/Z_{\bullet}(N_{\bullet})\bigr)\ \cong\ H_{r-1}\!\bigl(Z_{\bullet}(N_{\bullet})\bigr)\cong Z_{r-1}(N_{\bullet}).$
Remark 7.10 (Homology spheres).
For a normalized cycle $C\in Z_{n}(N_{\bullet})$, the generated simplicial subobject $\langle C\rangle\subseteq X_{\bullet}$ (the degeneracy-closure) provides an algebraic model of the $n$-sphere. If $C$ is primitive (e.g., a basis element $E_{m}$ over a PID $A$), $\langle C\rangle$ is isomorphic to the simplicial sphere $A[\Delta^{n}]/\operatorname{Sk}_{n-1}(A[\Delta^{n}])$. It satisfies $H_{n}(\langle C\rangle)\cong A$ and $H_{m}(\langle C\rangle)=0$ for $m\neq n$. (See Weibel [Weibel, Exercise 8.3.4].)
8 The Geometry of Generated Subobjects and Moduli Spaces
We analyze the isomorphism classes of the simplicial submodules $\langle T\rangle\subseteq X_{\bullet}(\vec{s};K)$ generated by a tensor $T\in X_{n}$. We assume the ground ring $K$ is an infinite field.
8.1 Realization Maps and Kernel Sequences
Let $C_{\bullet}:=K[\Delta^{n}]$ be the standard simplicial $K$-module representing the $n$-simplex. In degree $p$, $C_{p}=K[\Delta^{n}]_{p}$ is the free $K$-vector space with basis the set of morphisms $\Delta([p],[n])$. The dimension is $S_{p}:=\dim(C_{p})=\binom{n+p+1}{p+1}$. Let $\iota_{n}\in C_{n}$ be the generator corresponding to the identity map.
Definition 8.1 (Realization Map and Kernel Sequence).
For $T\in X_{n}(\vec{s};K)$, the realization map $f_{T}:C_{\bullet}\to X_{\bullet}(\vec{s};K)$ is the unique morphism of simplicial modules sending $\iota_{n}$ to $T$. The kernel sequence of $T$ is the simplicial submodule
$K(T)_{\bullet}:=\ker(f_{T})\subset C_{\bullet},$
defined degreewise by $K(T)_{p}:=\ker(f_{T,p})\subseteq C_{p}$.
The image is $\langle T\rangle_{\bullet}=\operatorname{im}(f_{T})$, and the realization map induces an isomorphism of simplicial $K$-modules
$C_{\bullet}/K(T)_{\bullet}\;\xrightarrow{\ \cong\ }\;\langle T\rangle_{\bullet}.$
The isomorphism class of $\langle T\rangle_{\bullet}$ is therefore determined by its kernel sequence $K(T)_{\bullet}$.
Remark 8.2 (Finite Generation).
Since $K[\Delta^{n}]$ is generated in degree $n$, any simplicial submodule is determined by its components up to degree $n$.
Proposition 8.3 (Quotients and generated submodules).
For each $T\in X_{n}(\vec{s};K)$, the realization map $f_{T}:C_{\bullet}\to X_{\bullet}(\vec{s};K)$ induces an isomorphism
$C_{\bullet}/K(T)_{\bullet}\xrightarrow{\ \cong\ }\langle T\rangle_{\bullet}.$
In particular, all invariants of the generated submodule $\langle T\rangle_{\bullet}$ can be computed from its kernel sequence $K(T)_{\bullet}\subseteq C_{\bullet}$.
Proof.
Since $K(T)_{\bullet}=\ker(f_{T})$, the universal property of quotients gives a simplicial map
$C_{\bullet}/K(T)_{\bullet}\longrightarrow\langle T\rangle_{\bullet}$
which is an isomorphism in each degree, because $K(T)_{p}=\ker(f_{T,p})$ and $\langle T\rangle_{p}=\operatorname{im}(f_{T,p})$. ∎
Remark 8.4.
For our purposes it suffices that each generated submodule $\langle T\rangle_{\bullet}$ arises as a quotient $C_{\bullet}/K(T)_{\bullet}$, and that all invariants we study (e.g. homology groups and the incidence conditions defining the moduli space) depend only on the kernel sequence $K(T)_{\bullet}\subseteq C_{\bullet}$. We make no assertion that isomorphic quotients have isomorphic kernels, nor that every isomorphism between quotients lifts to an automorphism of $C_{\bullet}$.
8.2 Index Collisions and the Realization Matrix
Let $R_{p}:=\dim(X_{p})$.
Definition 8.5 (Index collision map).
For each $p\geq 0$, the index collision map is
$\mathcal{I}_{p}:I_{p}\times\Delta([p],[n])\longrightarrow I_{n},\qquad\mathcal{I}_{p}(m,\alpha):=I_{\bullet}(\alpha)(m).$
Definition 8.6 (Realization matrix).
The realization matrix $M_{T,p}\in\operatorname{Mat}_{R_{p}\times S_{p}}(K)$ represents $f_{T,p}$. Its entry in row $m\in I_{p}$ and column $\alpha\in\Delta([p],[n])$ is
$M_{T,p}[m,\alpha]=(f_{T,p}(\alpha))(m)=T_{\mathcal{I}_{p}(m,\alpha)}.$
A collision in degree $p$ is a pair $(m,\alpha)\neq(m^{\prime},\alpha^{\prime})$ with $\mathcal{I}_{p}(m,\alpha)=\mathcal{I}_{p}(m^{\prime},\alpha^{\prime})$. Collisions force the corresponding entries of $M_{T,p}$ to agree for every $T$, and are the source of rank loss in the symbolic matrices $M_{T_{\mathrm{sym}},p}$.
8.3 The Moduli Map and Grassmannians
Let $\mathcal{V}=\{v_{u}\}_{u\in I_{n}}$ be indeterminates and $\mathbb{K}:=K(\mathcal{V})$ the function field. The symbolic tensor $T_{\text{sym}}$ has entries $v_{u}$.
Definition 8.7 (Generic Rank and Kernel Dimension).
The generic rank $R^{\prime}_{p}(\vec{s})$ is the rank of the symbolic matrix $M_{T_{\text{sym}},p}$ over $\mathbb{K}$. The generic kernel dimension is $K^{\prime}_{p}(\vec{s}):=S_{p}-R^{\prime}_{p}(\vec{s})$.
Definition 8.8 (Moduli Map).
The generic locus $\mathcal{U}\subset X_{n}(\vec{s};K)$ is the non-empty Zariski open set where $\operatorname{rank}(M_{T,p})=R^{\prime}_{p}(\vec{s})$ for all $p$.
The Grassmannian associated to the shape $\vec{s}$ is
$\operatorname{Gr}(\vec{s}):=\prod_{\begin{subarray}{c}0\leq p\leq n\\ K^{\prime}_{p}(\vec{s})>0\end{subarray}}\operatorname{Gr}\bigl(K^{\prime}_{p}(\vec{s}),K[\Delta^{n}]_{p}\bigr).$
By Remark 8.2, a simplicial submodule of $C_{\bullet}=K[\Delta^{n}]$ is determined by its components in degrees $0\leq p\leq n$, so recording the kernels $K(T)_{p}=\ker f_{T,p}$ in this range determines the entire kernel sequence $K(T)_{\bullet}$.
The Moduli Map $\Psi$ is the algebraic map:
$\Psi:\mathcal{U}\longrightarrow\operatorname{Gr}(\vec{s}),\quad T\mapsto(\ker f_{T,p})_{0\leq p\leq n}.$
The Moduli Space of Kernel Sequences $\mathcal{M}(\vec{s})$ is the image $\Psi(\mathcal{U})$.
8.4 Injectivity Analysis and Examples
Lemma 8.9.
The functor $I_{\bullet}:\Delta\to\mathbf{Set}$ preserves monomorphisms.
Proof.
Monomorphisms in $\Delta$ are generated by coface maps $\delta_{i}$. $I_{\bullet}(\delta_{i}^{p})=\Delta_{i}^{p}$ is a product of injections, hence injective. ∎
Theorem 8.10 (Injectivity and Dominance at $p=0$).
In the DSTM:
  1. 1.
    The map $\mathcal{I}_{0}$ is injective. $R^{\prime}_{0}=\min(R_{0},S_{0})$.
  2. 2.
    If $K^{\prime}_{0}>0$, the projection $\Psi_{0}:\mathcal{U}\to\operatorname{Gr}(K^{\prime}_{0},S_{0})$ is dominant.
Proof.
(1) $S_{0}=n+1$. Let $a_{0}$ be an axis where $n_{a_{0}}=n+1$. Then $M_{a_{0}}(0)=0$. For $m\in I_{0}$, $m_{a_{0}}=0$. Let $\alpha_{i}:[0]\to[n]$ be the vertex map $\alpha_{i}(0)=i$. The $a_{0}$-th coordinate of $\mathcal{I}_{0}(m,\alpha_{i})$ is $\alpha_{i}(m_{a_{0}})=i$. If $\mathcal{I}_{0}(m,\alpha_{i})=\mathcal{I}_{0}(m^{\prime},\alpha_{i^{\prime}})$, then $i=i^{\prime}$. By Lemma 8.9, $I_{\bullet}(\alpha_{i})$ is injective, so $m=m^{\prime}$.
(2) Since $\mathcal{I}_{0}$ is injective, the entries of $M_{T_{\text{sym}},0}$ are distinct variables. The map $L_{0}:X_{n}\to\operatorname{Mat}_{R_{0}\times S_{0}}(K)$ sending $T$ to $M_{T,0}$ is a projection onto these coordinates, hence surjective. The map from the space of matrices of maximal rank $R^{\prime}_{0}$ to the Grassmannian of their kernels is dominant. $\Psi_{0}$ is the composition of $L_{0}$ (restricted to $\mathcal{U}$) and this dominant map. ∎
Lemma 8.11 (Non-injectivity for $p\geq 1$).
For $\geq 1, p\mathcal{I}_{p}$ is not injective.
Proof.
Let $\alpha_{0}:[p]\to[n]$ be the constant map to 0. It factors as $\alpha_{0}=\iota_{0}\circ\pi_{0}$. Since $p\geq 1$, $|I_{p}|>|I_{0}|$. The map $I_{\bullet}(\pi_{0}):I_{p}\to I_{0}$ cannot be injective. There exist $m\neq m^{\prime}$ such that $I_{\bullet}(\pi_{0})(m)=I_{\bullet}(\pi_{0})(m^{\prime})$. Applying $I_{\bullet}(\iota_{0})$ yields $\mathcal{I}_{p}(m,\alpha_{0})=\mathcal{I}_{p}(m^{\prime},\alpha_{0})$. ∎
Example 8.12 (Rank Drop: Shape (2,2)).
$n=1,k=2$. $S=(2,3)$. $R=(1,4)$. $p=1$. $S_{1}=3,R_{1}=4$. The morphisms $\Delta([1],[1])$ are $\{\mathrm{id},(0,0),(1,1)\}$. $I_{1}=[1]^{2}$. The symbolic realization matrix $M_{T_{\text{sym}},1}$ is:
$M=\begin{pmatrix}v_{00}&v_{00}&v_{11}\\ v_{01}&v_{00}&v_{11}\\ v_{10}&v_{00}&v_{11}\\ v_{11}&v_{00}&v_{11}\end{pmatrix}$.
All four $3\times 3$ minors vanish identically over $\mathbb{K}$. The generic rank is $R^{\prime}_{1}=2$. $K^{\prime}_{1}=1$.
8.5 The Incidence Variety and Segre–Plücker Embedding
The image $\mathcal{M}(\vec{s})$ lies within the closed incidence subvariety $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})\subset\operatorname{Gr}(\vec{s})$ defined by the simplicial compatibility conditions.
Theorem 8.13 (Global Structure of the Moduli Space).
$\mathcal{M}(\vec{s})$ is a constructible set contained in $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$. The projection $\mathcal{M}(\vec{s})\to\operatorname{Gr}(K^{\prime}_{0}(\vec{s}),S_{0})$ is dominant (if $K^{\prime}_{0}(\vec{s})>0$). The moduli space $\mathcal{M}(\vec{s})$ is irreducible and unirational.
Proof.
$\mathcal{U}$ is an open subset of an affine space, hence irreducible. $\mathcal{M}(\vec{s})$ is the image of an irreducible variety under a rational map, hence it is constructible (Chevalley’s theorem), irreducible, and unirational. Dominance at $p=0$ follows from Theorem 8.10. ∎
Proposition 8.14 (Segre–Plücker linearity).
Embed $\operatorname{Gr}(\vec{s})$ into a projective space via the composition of the Plücker embeddings $\iota_{p}:\operatorname{Gr}(K^{\prime}_{p},S_{p})\hookrightarrow\mathbb{P}(\Lambda^{K^{\prime}_{p}}K[\Delta^{n}]_{p})$ and the Segre embedding $\Sigma$. The image $\Sigma(\mathbf{Gr}^{\mathrm{simp}}(\vec{s}))$ is the intersection of $\Sigma(\operatorname{Gr}(\vec{s}))$ with a linear subspace of the ambient projective space.
Proof.
We analyze the condition $d_{i}(K_{p})\subseteq K_{p-1}$. Let $C_{q}=K[\Delta^{n}]_{q}$. Let $\omega_{p}$ and $\omega_{p-1}$ be the Plücker coordinates representing $K_{p}$ and $K_{p-1}$. The condition is equivalent to $\Lambda^{K^{\prime}_{p}}d_{i}(\omega_{p})\wedge\eta_{p-1}=0$, where $\eta_{p-1}$ is a volume element of a complement to $K_{p-1}$. Under the Segre embedding, these bilinear equations in the Plücker coordinates become linear equations in the homogeneous coordinates of the tensor product space. ∎
Proposition 8.15 (Determinantal structure of the incidence variety).
Let $C_{p}:=K[\Delta^{n}]_{p}$. On $\operatorname{Gr}(\vec{s})$, let $\mathcal{S}_{p}\subset C_{p}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}$ denote the pullback of the tautological rank-$K^{\prime}_{p}(\vec{s})$ subbundle on $\operatorname{Gr}(K^{\prime}_{p}(\vec{s}),C_{p})$ along the projection $\operatorname{Gr}(\vec{s})\to\operatorname{Gr}(K^{\prime}_{p}(\vec{s}),C_{p})$, and let $\mathcal{S}_{p-1}$ be defined similarly in degree p1p-1.
For each pair $(p,i)$ with $K^{\prime}_{p}(\vec{s})>0$, the face map $d_{i}:C_{p}\to C_{p-1}$ induces a morphism of vector bundles
$\varphi_{p,i}:\mathcal{S}_{p}\longrightarrow\bigl(C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}\bigr)/\,\mathcal{S}_{p-1}.$
Then the incidence variety $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$ is the scheme-theoretic intersection
$\mathbf{Gr}^{\mathrm{simp}}(\vec{s})=\bigcap_{p,i}\{\,\varphi_{p,i}=0\,\},$
i.e. it is the intersection of the rank-0 determinantal loci of the bundle maps $\varphi_{p,i}$. In particular, $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$ is a determinantal subvariety of $\operatorname{Gr}(\vec{s})$.
Proof.
A point of $\operatorname{Gr}(\vec{s})$ is a tuple of subspaces $(K_{p})_{p}$ with $K_{p}\subset C_{p}$ and $\dim K_{p}=K^{\prime}_{p}(\vec{s})$ whenever $K^{\prime}_{p}(\vec{s})>0$. By construction of the tautological subbundle, the fiber $(\mathcal{S}_{p})_{x}$ over a point $x$ corresponding to $(K_{p})_{p}$ is exactly $K_{p}$, and similarly $(\mathcal{S}_{p-1})_{x}=K_{p-1}$.
The fixed linear map $d_{i}:C_{p}\to C_{p-1}$ induces a morphism of vector bundles
$d_{i}\otimes\mathrm{id}:C_{p}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}\longrightarrow C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}.$
Composing with the quotient map $C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})}\twoheadrightarrow(C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})})/\mathcal{S}_{p-1}$ and restricting to $\mathcal{S}_{p}\subset C_{p}\otimes\mathcal{O}$ gives a bundle map
$\varphi_{p,i}:\mathcal{S}_{p}\to(C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})})/\mathcal{S}_{p-1}.$
At a point $x=(K_{p})_{p}$, the fiber map $(\varphi_{p,i})_{x}:K_{p}\to C_{p-1}/K_{p-1}$ is the composition
$K_{p}\hookrightarrow C_{p}\xrightarrow{d_{i}}C_{p-1}\twoheadrightarrow C_{p-1}/K_{p-1}.$
Thus $(\varphi_{p,i})_{x}=0$ if and only if $d_{i}(K_{p})\subseteq K_{p-1}$. By definition, $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$ is the locus of tuples satisfying these inclusions for all $(p,i)$.
In local trivializations, $\varphi_{p,i}$ is represented by a matrix of regular functions, and the condition $(\varphi_{p,i})_{x}=0$ corresponds to the vanishing of all entries (all $1\times 1$ minors). Globally, this condition defines the zero locus of the section $\varphi_{p,i}$ of the bundle $\mathcal{H}om(\mathcal{S}_{p},\mathcal{Q}_{p-1})$, where we define the quotient bundle $\mathcal{Q}_{p-1}:=(C_{p-1}\otimes\mathcal{O}_{\operatorname{Gr}(\vec{s})})/\mathcal{S}_{p-1}$. In terms of homogeneous coordinates, this condition is bilinear in the Plücker coordinates of the factors (and thus linear in the Segre coordinates), confirming that $\mathbf{Gr}^{\mathrm{simp}}(\vec{s})$ is a determinantal subvariety. ∎
8.6 Homology of Generated Subobjects
Fix $T\in X_{n}(\vec{s};K)$ and write $C_{\bullet}:=K[\Delta^{n}]$ and $K_{\bullet}:=K(T)_{\bullet}=\ker(f_{T})\subseteq C_{\bullet}$. We analyze $H_{\bullet}(\langle T\rangle_{\bullet})$ using the short exact sequence
$0\longrightarrow K_{\bullet}\longrightarrow C_{\bullet}\longrightarrow\langle T\rangle_{\bullet}\longrightarrow 0.$
Recall that $H_{p}(C_{\bullet})=0$ for $p>0$ and $H_{0}(C_{\bullet})\cong K$.
Proposition 8.16 (Homology of the generated subobject).
Let $B_{p}(C_{\bullet})$ and $B_{p}(K_{\bullet})$ denote the boundary subspaces in $C_{\bullet}$ and $K_{\bullet}$, respectively. Then:
  1. 1.
    For $p\geq 2$, there is a natural isomorphism
    $H_{p}(\langle T\rangle_{\bullet})\cong H_{p-1}(K_{\bullet}).$
  2. 2.
    $H_{1}(\langle T\rangle_{\bullet})\cong\bigl(K_{0}\cap B_{0}(C_{\bullet})\bigr)/B_{0}(K_{\bullet})$.
  3. 3.
    $H_{0}(\langle T\rangle_{\bullet})\cong C_{0}/\bigl(K_{0}+B_{0}(C_{\bullet})\bigr)$.
Proof.
Applying homology to
$0\to K_{\bullet}\to C_{\bullet}\to\langle T\rangle_{\bullet}\to 0$
gives a long exact sequence. For $p\geq 2$ we have $H_{p}(C_{\bullet})=H_{p-1}(C_{\bullet})=0$, so the connecting homomorphism induces an isomorphism $H_{p}(\langle T\rangle_{\bullet})\cong H_{p-1}(K_{\bullet})$, which proves (1).
For $p=1$ and $p=0$, the long exact sequence reads
$0\to H_{1}(\langle T\rangle_{\bullet})\to H_{0}(K_{\bullet})\to H_{0}(C_{\bullet})\to H_{0}(\langle T\rangle_{\bullet})\to 0,$
and the standard identification of $H_{0}$ as cycles modulo boundaries in degree 0 yields the descriptions in (2) and (3). ∎
Corollary 8.17 (Generic connectivity).
If $K^{\prime}_{0}(\vec{s})>0$, then for a generic tensor $T\in\mathcal{U}$, $H_{0}(\langle T\rangle_{\bullet})=0$.
Proof.
$B_{0}(C_{\bullet})$ is a hyperplane in $C_{0}$. By Proposition 8.16, we have $H_{0}(\langle T\rangle_{\bullet})\neq 0$ if and only if $K_{0}(T)\subseteq B_{0}(C_{\bullet})$. This condition cuts out a proper closed subvariety of $\operatorname{Gr}(K^{\prime}_{0}(\vec{s}),S_{0})$. Since $\Psi_{0}$ is dominant (Theorem 8.10), the generic kernel $K_{0}(T)$ avoids this locus. ∎
8.7 Example: Shape $(3,3)$
Let $\vec{s}=(3,3)$, so $k=2$ and $n=2$. Then $X_{2}$ consists of $3\times 3$ matrices. We write $T=(v_{ij})_{0\leq i,j\leq 2}$. The dimensions are
$S=(3,6,10),\qquad R=(1,4,9).$
We fix the following bases:
  • 1.
    For $X_{2}$, the basis $I_{2}=[2]^{2}$ in lexicographic order:
    (0,0),(0,1),(0,2),(1,0),(1,1),(1,2),(2,0),(2,1),(2,2).(0,0),(0,1),(0,2),(1,0),(1,1),(1,2),(2,0),(2,1),(2,2).
  • 2.
    For $K[\Delta^{2}]_{1}$, the basis of monotone maps $1]\to[2]$ ordered as
    (0,0),(0,1),(0,2),(1,1),(1,2),(2,2).(0,0),\ (0,1),\ (0,2),\ (1,1),\ (1,2),\ (2,2).
  • 3.
    For $K[\Delta^{2}]_{2}$, the basis of monotone maps $2]\to[2]$ ordered as the ten non-decreasing triples $(a_{0},a_{1},a_{2})$ with $0\leq a_{0}\leq a_{1}\leq a_{2}\leq 2$:
    (0,0,0),(0,0,1),(0,0,2),(0,1,1),(0,1,2),(0,2,2),(1,1,1),(1,1,2),(1,2,2),(2,2,2).(0,0,0),(0,0,1),(0,0,2),(0,1,1),(0,1,2),(0,2,2),(1,1,1),(1,1,2),(1,2,2),(2,2,2).
Degree $p=0$.
Here $R_{0}=1$ and $S_{0}=3$. The realization matrix is
$M_{T,0}=\begin{pmatrix}v_{00}&v_{11}&v_{22}\end{pmatrix}$.
For generic $T$, its rank is $1$, so $K^{\prime}_{0}=2$. The kernel is the $2$-dimensional subspace of $K^{3}$ of solutions to
$v_{00}x_{0}+v_{11}x_{1}+v_{22}x_{2}=0.$
On the open chart $v_{00}\neq 0$, it is spanned by
$(-v_{11}/v_{00},\,1,\,0),\qquad(-v_{22}/v_{00},\,0,\,1).$
Thus $K_{0}(T)$ depends only on the projective class $v_{00}:v_{11}:v_{22}]\in\mathbb{P}^{2}$.
Degree $p=1$.
Here $R_{1}=4$ and $S_{1}=6$. With the bases above, the realization matrix is
$M_{T,1}=\begin{pmatrix}v_{00}&v_{00}&v_{00}&v_{11}&v_{11}&v_{22}\\ v_{00}&v_{01}&v_{02}&v_{11}&v_{12}&v_{22}\\ v_{00}&v_{10}&v_{20}&v_{11}&v_{21}&v_{22}\\ v_{00}&v_{11}&v_{22}&v_{11}&v_{22}&v_{22}\end{pmatrix}$.
Two obvious kernel vectors can be written purely in terms of the diagonal entries. Set
[…]
A direct calculation shows
$M_{T,1}\,c^{(1)}_{1}=0,\qquad M_{T,1}\,c^{(1)}_{2}=0$
as identities in the polynomial ring $K[v_{ij}]$ (on the open set $v_{00}\neq 0$). Hence $K^{\prime}_{1}\geq 2$.
On the other hand, taking e.g.
$T=\begin{pmatrix}1&1&1\\ 2&2&1\\ 2&2&3\end{pmatrix}$,
one obtains
$M_{T,1}=\begin{pmatrix}1&1&1&2&2&3\\ 1&1&1&2&1&3\\ 1&2&2&2&2&3\\ 1&2&3&2&3&3\end{pmatrix}$,
which has rank $4$ (one checks that some $4\times 4$ minor is non-zero). Thus the generic rank is $R^{\prime}_{1}=4$, and therefore $K^{\prime}_{1}=S_{1}-R^{\prime}_{1}=2$.
In particular, for generic $T$, the kernel $K_{1}(T)=\ker f_{T,1}$ is spanned by $c^{(1)}_{1}$ and $c^{(1)}_{2}$, so it depends only on $(v_{00},v_{11},v_{22})$.
Degree $p=2$.
Here $R_{2}=9$ and $S_{2}=10$. With the bases chosen above, the symbolic realization matrix $M_{T,2}$ is
$M_{T,2}=\begin{pmatrix}v_{00}&v_{00}&v_{00}&v_{11}&v_{11}&v_{22}&v_{11}&v_{11}&v_{22}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{01}&v_{01}&v_{02}&v_{11}&v_{11}&v_{12}&v_{22}\\ v_{00}&v_{00}&v_{00}&v_{10}&v_{10}&v_{20}&v_{11}&v_{11}&v_{21}&v_{22}\\ v_{00}&v_{00}&v_{01}&v_{11}&v_{12}&v_{22}&v_{11}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{01}&v_{02}&v_{11}&v_{12}&v_{22}&v_{11}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{02}&v_{02}&v_{12}&v_{12}&v_{22}&v_{12}&v_{12}&v_{22}&v_{22}\\ v_{00}&v_{00}&v_{10}&v_{11}&v_{21}&v_{22}&v_{11}&v_{21}&v_{22}&v_{22}\\ v_{00}&v_{10}&v_{20}&v_{11}&v_{21}&v_{22}&v_{11}&v_{21}&v_{22}&v_{22}\\ v_{00}&v_{20}&v_{20}&v_{21}&v_{21}&v_{22}&v_{21}&v_{21}&v_{22}&v_{22}\end{pmatrix}$.
Again there are two explicit kernel vectors depending only on the diagonal. Set
[…]
A direct substitution into the matrix above shows
$M_{T,2}\,c^{(2)}_{1}=0,\qquad M_{T,2}\,c^{(2)}_{2}=0$
as polynomial identities (again on the open chart $v_{00}\neq 0$). Thus $K^{\prime}_{2}\geq 2$ and $R^{\prime}_{2}\leq 10-2=8$.
To see that the generic rank is exactly $8$, we exhibit a single tensor for which $M_{T,2}$ has rank $8$. Take for example
$T=\begin{pmatrix}1&1&1\\ 2&2&1\\ 2&2&3\end{pmatrix}$.
Substituting these values into $M_{T,2}$ gives a $9\times 10$ numerical matrix with rank $8$. Since rank is upper semicontinuous in families, this shows that the generic rank is $R^{\prime}_{2}=8$, and therefore $K^{\prime}_{2}=S_{2}-R^{\prime}_{2}=2$.
Moreover, the two kernel generators $c^{(2)}_{1},c^{(2)}_{2}$ depend only on $v_{00},v_{11},v_{22}$, not on the off-diagonal entries.
Collecting the three degrees, we have:
Proposition 8.18 (Generic kernel dimensions for shape $(3,3)$).
For $\vec{s}=(3,3)$, the generic kernel dimensions are
$K^{\prime}(\vec{s})=(K^{\prime}_{0},K^{\prime}_{1},K^{\prime}_{2})=(2,2,2).$
On the open chart $v_{00}\neq 0$, the kernels $K_{p}(T)$ are generated by vectors whose coordinates are rational functions of $v_{11}/v_{00}$ and $v_{22}/v_{00}$ only.
Corollary 8.19 (Moduli space for shape $(3,3)$).
For $\vec{s}=(3,3)$, the moduli space $\mathcal{M}(\vec{s})$ is a surface of dimension $2$, birational to $\mathbb{P}^{2}$.
Proof.
The generic kernel vector in degree 0 is the hyperplane
$K_{0}(T)=\{(x_{0},x_{1},x_{2})\mid v_{00}x_{0}+v_{11}x_{1}+v_{22}x_{2}=0\}\subset K^{3}.$
Thus the point $K_{0}(T)\in\operatorname{Gr}(2,3)$ determines the projective class $v_{00}:v_{11}:v_{22}]\in\mathbb{P}^{2}$, and hence determines the ratios $v_{11}/v_{00}$ and $v_{22}/v_{00}$ on the chart $v_{00}\neq 0$. By the explicit formulas above, $K_{1}(T)$ and $K_{2}(T)$ are then uniquely determined. Therefore, on this chart, the moduli map $\Psi$ factors through the projection
$X_{2}\dashrightarrow\mathbb{P}^{2},\qquad T\longmapsto[v_{00}:v_{11}:v_{22}],$
and the induced map $\mathbb{P}^{2}\dashrightarrow\mathcal{M}(\vec{s})$ is birational onto its image. Hence $\mathcal{M}(\vec{s})$ has dimension $2$ and is birational to $\mathbb{P}^{2}$. ∎
Appendix A Contractibility of the Diagonal Simplicial Module
We prove that the diagonal simplicial $A$-module $X_{\bullet}(\vec{s};A)$ is acyclic by constructing an explicit chain contraction.
A.1 The Chain Complex and Homotopy Operator
The DSTM $X_{p}(\vec{s};A)$ is the $A$-module of functions $T:I_{p}\to A$. The index bounds are $M_{a}(p)=n_{a}-1-n+p$, where $n=\min_{a}(n_{a})-1$.
Remark A.1 (Index Bounds at $p=0$).
By the definition of $n$, there exists at least one index $a_{0}$ such that $n_{a_{0}}-1=n$. For this index, $M_{a_{0}}(0)=0$. Consequently, for any index $m\in I_{0}$, the coordinate $m_{a_{0}}$ must be 0.
Let $(X_{\bullet},\partial_{\bullet})$ be the associated unaugmented chain complex. We set $X_{p}=0$ for $p<0$, and $\partial_{0}=0$.
Definition A.2 (Shift-and-Truncate Homotopy $H$).
We define $H_{p}:X_{p}\to X_{p+1}$ for $p\geq-1$. Set $H_{-1}:=0$. For $p\geq 0$, $T\in X_{p}$, and $m\in I_{p+1}$:
$H_{p}(T)(m):=\begin{cases}0&\text{if }\exists a:m_{a}=0,\\ T(m-\vec{1})&\text{if }\forall a:m_{a}>0.\end{cases}$
Lemma A.3.
The operator $H$ satisfies the following identities:
  1. 1.
    For all $p\geq 0$, $d_{0}^{p+1}H_{p}=\mathrm{id}_{X_{p}}$.
  2. 2.
    For all $p\geq 1$ and all $i>0$, $d_{i}^{p+1}H_{p}=H_{p-1}d_{i-1}^{p}$.
Proof.
Let $T\in X_{p}$ and $m\in I_{p}$.
(1) $d_{0}H(T)(m)=H(T)(\Delta_{0}(m))=T(m+\vec{1}-\vec{1})=T(m)$.
(2) Let $i>0$.
Case $p>0$: If some coordinate $m_{a}=0$, then $\Delta_{i}(m)$ also has a zero coordinate (since $i>0$ and $\delta_{i}(0)=0$), so $H_{p}(T)(\Delta_{i}(m))=0$ and $H_{p-1}(d_{i-1}T)(m)=0$ by definition of $H$. Thus $d_{i}H_{p}(T)(m)=H_{p-1}d_{i-1}(T)(m)=0$.
If all coordinates $m_{a}>0$, then
$d_{i}^{p+1}H_{p}(T)(m)=H_{p}(T)(\Delta_{i}^{p}(m))=T(\Delta_{i}^{p}(m)-\vec{1}),$
while
$H_{p-1}d_{i-1}^{p}(T)(m)=d_{i-1}^{p}(T)(m-\vec{1})=T(\Delta_{i-1}^{p}(m-\vec{1})).$
The coface maps satisfy the coordinate identity
$\delta_{i}(x)-1=\delta_{i-1}(x-1)\qquad(x>0),$
hence $\Delta_{i}^{p}(m)-\vec{1}=\Delta_{i-1}^{p}(m-\vec{1})$ when all $m_{a}>0$, and the two expressions agree.
Case $p=0$: We verify $d_{1}^{1}H_{0}=0$.
Let $m\in I_{0}$. By Remark A.1, there exists $a_{0}$ such that $m_{a_{0}}=0$. $d_{1}H_{0}(T)(m)=H_{0}(T)(\Delta_{1}(m))$. The $a_{0}$-th coordinate of $\Delta_{1}(m)$ is $\delta_{1}(m_{a_{0}})=\delta_{1}(0)$. Since $i=1>0$, $\delta_{1}(0)=0$. By the definition of $H$, since $\Delta_{1}(m)$ has a zero coordinate, $H_{0}(T)(\Delta_{1}(m))=0$. Thus $d_{1}^{1}H_{0}=0$. ∎
Theorem A.4.
The diagonal simplicial module $X_{\bullet}(\vec{s};A)$ is contractible.
Proof.
We verify the chain contraction identity $\partial_{p+1}H_{p}+H_{p-1}\partial_{p}=\mathrm{id}$. Using Lemma A.3 (valid for all $p\geq 0$):
$\partial_{p+1}H_{p}$ $=d_{0}H_{p}+\sum_{i=1}^{p+1}(-1)^{i}d_{i}H_{p}$
$=\mathrm{id}+\sum_{i=1}^{p+1}(-1)^{i}H_{p-1}d_{i-1}^{p}.$
Re-indexing the sum (j=i1j=i-1):
$\sum_{i=1}^{p+1}(-1)^{i}H_{p-1}d_{i-1}^{p}$ $=\sum_{j=0}^{p}(-1)^{j+1}H_{p-1}d_{j}^{p}=-\,H_{p-1}\partial_{p}.$
(This holds even for $p=0$, as $\partial_{0}=0$ and $H_{-1}=0$). Substituting back, we obtain
$\partial_{p+1}H_{p}+H_{p-1}\partial_{p}\;=\;\mathrm{id}_{X_{p}},$
so $H$ is a chain contraction of $X_{\bullet}(\vec{s};A)$. ∎
A.2 Equivariance Properties of the Homotopy
Recall the action of the stabilizer group $\operatorname{Stab}(\vec{s})$ on the diagonal simplicial module $X_{\bullet}(\vec{s};A)$ defined in Section 2. The face and degeneracy maps commute with this action (Lemma 2.4).
Lemma A.5 (Equivariance under $\operatorname{Stab}(\vec{s})$).
For every $p\geq-1$ and $\sigma\in\operatorname{Stab}(\vec{s})$, the shift-and-truncate homotopy $H_{p}$ is equivariant with respect to the action of $\operatorname{Stab}(\vec{s})$:
$H_{p}(\sigma\cdot T)=\sigma\cdot H_{p}(T).$
Proof.
Let $T\in X_{p}$, $\sigma\in\operatorname{Stab}(\vec{s})$, and $m\in I_{p+1}$. We compare $H_{p}(\sigma\cdot T)(m)$ and $(\sigma\cdot H_{p}(T))(m)$. By definition of the action, $(\sigma\cdot H_{p}(T))(m)=H_{p}(T)(\sigma^{-1}m)$.
The definition of $H_{p}$ depends on the presence of zero coordinates. Since permutation only changes the positions of the coordinates, the condition $\exists a:m_{a}=0$” is equivalent to $\exists b:(\sigma^{-1}m)_{b}=0$”.
Case 1: $\exists a:m_{a}=0$. In this case, $H_{p}(\sigma\cdot T)(m)=0$. Since $\sigma^{-1}m$ also contains a zero coordinate, $H_{p}(T)(\sigma^{-1}m)=0$.
Case 2: $\forall a:m_{a}>0$. In this case, $\sigma^{-1}m$ is also strictly positive.
$H_{p}(\sigma\cdot T)(m)$ $=(\sigma\cdot T)(m-\vec{1})$
$=T(\sigma^{-1}(m-\vec{1})).$
On the other hand,
$(\sigma\cdot H_{p}(T))(m)$ $=H_{p}(T)(\sigma^{-1}m)$
$=T((\sigma^{-1}m)-\vec{1}).$
The equality holds because the shift operation commutes with permutation: $\sigma^{-1}(m-\vec{1})=(\sigma^{-1}m)-\vec{1}$. Therefore, $H_{p}$ is equivariant under the action of $\operatorname{Stab}(\vec{s})$. ∎
A.3 Shifted Depth Filtration and Spectral Sequence
We introduce a filtration on the complex $(X_{\bullet},\partial)$ that is compatible with the differential and the homotopy operator $H$, leading to the collapse of the associated spectral sequence.
Definition A.6 (Shifted Depth Filtration).
For $p\geq 0$ and $t\in\mathbb{Z}$, define a decreasing filtration $\mathcal{F}^{\bullet}X_{p}$. Write $\min(m):=\min_{a}m_{a}$ for $m\in I_{p}$. Set
$\mathcal{F}^{t}X_{p}\;:=\;\{\,T\in X_{p}\mid T(m)=0\ \text{unless }\min(m)\geq t+p\,\}.$
This filtration is bounded: $\mathcal{F}^{-p}X_{p}=X_{p}$ (since $\min(m)\geq 0$). Furthermore, since $\min(m)\leq\min_{a}M_{a}(p)$ for any $m\in I_{p}$, we have $\mathcal{F}^{t}X_{p}=0$ if $t+p>\min_{a}M_{a}(p)$.
Lemma A.7 (Behavior of $\partial$ and $H$ w.r.t. $\mathcal{F}^{\bullet}$).
The differential $\partial$ and the homotopy $H$ preserve the shifted depth filtration (filtration degree 0). For all $p\geq 0$ and $t\in\mathbb{Z}$:
  1. 1.
    $d_{i}^{p}(\mathcal{F}^{t}X_{p})\subseteq\mathcal{F}^{t}X_{p-1}$ for all $i\in\{0,\dots,p\}$.
  2. 2.
    $H_{p}(\mathcal{F}^{t}X_{p})\subseteq\mathcal{F}^{t}X_{p+1}$.
Proof.
(1) Let $T\in\mathcal{F}^{t}X_{p}$. We want to show $d_{i}^{p}(T)\in\mathcal{F}^{t}X_{p-1}$. Let $m\in I_{p-1}$. We must show that if $\min(m)<t+p-1$, then $d_{i}^{p}(T)(m)=0$. $d_{i}^{p}(T)(m)=T(\Delta_{i}^{p}(m))$. The coface maps satisfy $\delta_{i}(x)\leq x+1$ for all $i,x$. Thus $\min(\Delta_{i}^{p}(m))\leq\min(m)+1$. If $\min(m)<t+p-1$, then $\min(\Delta_{i}^{p}(m))<t+p$. Since $T\in\mathcal{F}^{t}X_{p}$, $T(\Delta_{i}^{p}(m))=0$. Thus $d_{i}^{p}(T)\in\mathcal{F}^{t}X_{p-1}$.
(2) Let $T\in\mathcal{F}^{t}X_{p}$. We want to show $H_{p}(T)\in\mathcal{F}^{t}X_{p+1}$. Let $m\in I_{p+1}$. We must show that if $\min(m)<t+p+1$, then $H_{p}(T)(m)=0$.
If $\min(m)=0$, $H_{p}(T)(m)=0$ by definition. If $\min(m)>0$, $H_{p}(T)(m)=T(m-\vec{1})$. If $\min(m)<t+p+1$, then $\min(m-\vec{1})=\min(m)-1<t+p$. Since $T\in\mathcal{F}^{t}X_{p}$, $T(m-\vec{1})=0$. Thus $H_{p}(T)\in\mathcal{F}^{t}X_{p+1}$. ∎
Proposition A.8 (Spectral sequence collapses at $E_{1}$).
Let $\mathcal{F}^{\bullet}$ be the shifted depth filtration. The spectral sequence of the filtered complex $(X_{\bullet},\partial,\mathcal{F}^{\bullet})$ satisfies
$E_{1}^{t,q}=0\qquad\text{for all }t,q.$
Hence it collapses at the $E_{1}$–-page.
Proof.
By Lemma A.7, $(X_{\bullet},\partial,\mathcal{F}^{\bullet})$ is a filtered complex, as $\partial$ has filtration degree 0. Let $E_{0}=\operatorname{gr}_{\mathcal{F}}(X_{\bullet})$ be the associated graded complex, with differential $d^{0}$ induced by $\partial$.
By Lemma A.7, the homotopy operator $H$ also has filtration degree 0. It induces a map $[H]$ on $E_{0}$. The chain contraction identity $\partial H+H\partial=\mathrm{id}$ holds in the filtered complex. Since all operators involved preserve the filtration, the identity descends to the associated graded complex:
$d^{0}[H]+[H]d^{0}=[\mathrm{id}]=\mathrm{id}_{E_{0}}.$
This shows that the complex $(E_{0},d^{0})$ is contractible via the homotopy $[H]$. Therefore, its homology is trivial. The $E_{1}$ term is defined as the homology of $(E_{0},d^{0})$. We conclude $E_{1}=H(E_{0})=0$. ∎